### v2

# Training with Pseudo-Code for Instruction Following

Abstract
Despite rapid advances in the capabilities of Large Language Models (LLMs), they continue to struggle with following relatively simple and unambiguous instructions, particularly when compositional structure is involved. Recent work suggests that models may follow instructions more effectively when they are expressed in pseudo-code rather than natural language (Mishra et al., 2023; Puerto et al., 2024). However, writing pseudo-code programs can be tedious, and relying on few-shot demonstrations or inference-time code prompting is often unnatural for non-expert users of LLMs. To overcome these limitations, we propose a training-time approach that fine-tunes LLMs using instruction-tuning data augmented with pseudo-code representations of natural-language instructions paired with final responses. We evaluate our method on 12 publicly available benchmarks spanning instruction-following, mathematical reasoning, and commonsense reasoning, across six base models. Our results show that models trained with pseudo-code follow instructions more reliably, achieving relative gains of 8–21% on instruction-following benchmarks, while largely preserving—and in some cases improving—performance on mathematical and commonsense reasoning tasks, with an average gain of up to 30% across all evaluated benchmarks.
Machine Learning, ICML
1 Introduction
Despite the rapid progress in the capabilities of Large Language Models (LLMs), they continue to face challenges in effectively processing and executing instructions across diverse tasks. These challenges manifest in several ways: difficulty with compositional instructions (Wen et al., 2024; Jiang et al., 2024b; Murthy et al., 2025), susceptibility to distractions (Hwang et al., 2025), and struggles with adhering to specified output formats (Xia et al., 2024). To address these limitations, significant research effort has focused on improving how models understand and respond to instructions. This includes enforcing constraints during decoding (Banerjee et al., 2025), prompting strategies that elicit structured thinking and planning behavior (Xia et al., 2025), test-time scaling (Snell et al., 2024; Bansal et al., 2025), multi-step agentic flows (Yao et al., 2023b; Mitra et al., 2024), and more. Additionally, training-time interventions have emerged as an active area of work, including the design of training objectives (Pham et al., 2024; Shao et al., 2024; Dong et al., 2025; DeepSeek-AI et al., 2025), incorporation of chain-of-thought reasoning (Longpre et al., 2023a; Mitra et al., 2024), and, crucially, careful curation of training data (Lou et al., 2024; Sun et al., 2024; Qi et al., 2024; He et al., 2024).
Figure 1: Sketch of the data conversion pipeline along with an example instruction and corresponding pseudo-code program. The prompt used for generating pseudo-code is shown in Appendix Figure 7 and the prompt used for pseudo-code Repair is presented in Appendix Figure 30.
In this paper, we draw inspiration from recent work showing that large language models often perform better when instructions are expressed in pseudo-code rather than natural language (Mishra et al., 2023; Li et al., 2024a; Puerto et al., 2024). Compared to natural language, pseudo-code provides a more explicit representation of task structure, including ordering, control flow, and execution constraints. Mishra et al. (2023) further show that code-specific elements—such as comments, docstrings, and structural cues, can improve model performance. By clarifying task structure and constraints, these elements yield instructions that are clearer, more concise, and less ambiguous. As a result, both directly prompting models with pseudo-code (Mishra et al., 2023) and using pseudo-code as a thinking aid during inference (Puerto et al., 2024) have been shown to improve performance across a wide range of natural language processing tasks, including text classification, question answering, question generation, and commonsense reasoning.
However, relying on pseudo-code at inference time introduces practical limitations. Writing pseudo-code programs can be tedious and unintuitive for non-expert users, and few-shot demonstrations that incorporate code representations are often brittle. Moreover, approaches such as Chain-of-Code (Li et al., 2024a) depend on external code execution or emulation to produce intermediate outputs, complicating deployment.
To address these limitations, we propose a training-time alternative. We fine-tune base LLMs using instruction-tuning data that pairs natural-language instructions with corresponding pseudo-code representations and final answers. Given a natural-language prompt, the model is trained to first re-express the instruction in pseudo-code and then generate the answer (see the sample program in Figure 1). From an end-user’s perspective, interaction with the resulting model is identical to standard natural-language prompting, making our approach a seamless drop-in replacement for existing instruction-tuned LLMs.
Training models to re-express instructions in pseudo-code before answering a user’s request is similar in spirit to approaches that train models to generate hidden chain-of-thought (CoT) or “thinking” steps (Hayati et al., 2025). However, in this work we do not enrich pseudo-code instructions (programs) with such “thinking” steps; instead, we re-express instructions solely to isolate the impact of pseudo-code–based instruction training on LLM performance.  [^1] Note: Our training data includes some tasks where CoT is expressed in final responses (but not in programs).
Through experiments on publicly available benchmarks, we find that models trained with pseudo-code–augmented instruction tuning follow instructions more reliably and robustly than their natural-language–trained counterparts. Moreover, these gains do not come at the expense of general capability: pseudo-code–trained models retain—and in some cases improve—their performance on tasks other tasks such as those involving mathematical and commonsense reasoning.
Contributions: We propose a training-time approach for improving instruction-following in large language models by learning a structured pseudo-code intermediate representation of natural-language instructions. Specifically, (1) we train models to re-express instructions in pseudo-code before answering, improving execution of compositional, constrained, and nested instructions without inference-time prompting or external code execution; (2) we introduce a scalable generate–evaluate–repair pipeline that automatically augments instruction-tuning data with high-quality pseudo-code supervision; and (3) through experiments on six base models and twelve benchmarks, we show performance improvements of (8–21% relative) on instruction-following tasks while largely preserving or improving performance on mathematical, commonsense, and code-related tasks, and outperforming inference-time code prompting baselines.
2 Data Construction
Data Generation Pipeline: As mentioned previously, we enhance existing instruction-tuning datasets by re-expressing instructions in pseudo-code as part of the output. Since manually crafting pseudo-code programs for instruction-tuning is prohibitively expensive, we develop a multi-step data conversion pipeline to create pseudo-code instruction tuning data. Our pipeline consists of three stages: (i) Generate: Use a powerful LLM with a 1-shot prompt to generate pseudo-code programs for a given instruction. We manually write 5 pseudo-code instructions for each of the datasets. We randomly select $1$ of them for the 1-shot prompt. (ii) Evaluate: Similar to (Mishra et al., 2023), use only the generated pseudo-code program to run inference and get the output of the instruction. We compare and score the generated output against the ground-truth output, (iii) Repair: For the cases when the inference using the pseudo-code program does not result in the correct ground-truth instruction output, but it does when inference is run using the original natural language instruction, we prompt an LLM-judge to give feedback to improve the pseudo-code program. We summarize the pipeline in Figure 1.
In order to generate pseudo-code programs, the pipeline uses hand-crafted demonstrations of pseudo-code instructions from (Mishra et al., 2023) with Mixtral-8x7B-Instruct-v0.1222https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1 as the teacher model. Prompts for different stages of this pipeline are available in Appendix Figures 7 and 30. All our pseudo code programs also include a statement mimicking the “invocation” of the program (See Line $13$ in the Pseudo-code program in Figure 1). An instruction tuning instance consists of the original NL-instruction as the input along with the generated pseudo-code program and the original NL-instruction output text (separated by custom [PSEUDOCODE] markers), as the output of the instruction instance (see example instances in Appendix Section D).
Dataset #Samples Sampling Proportion Code Alpaca (Chaudhary, 2023) 20016 0.15 Tulu v2 Science (Rafailov et al., 2024) 7468 0.0375 SNI v2 (Wang et al., 2022) 85187 0.325 WizardLM Evol Instruct (Xu et al., 2024) 26180 0.15 GPT4-Alpaca (Peng et al., 2023) 46149 0.15 CoT (Longpre et al., 2023b) 98315 0.1875
Table 1: Statistics of our Training Mixture from the TuluV2 dataset (Ivison et al., 2023).
Training Mixture: For our experiments, we use datasets from the Tulu v2 mixture (Ivison et al., 2023). We obtain the full set of instances for Code Alpaca (Chaudhary, 2023) and GPT4-Alpaca (Peng et al., 2023) instead of the subset of instances present in Tuluv2 mixture. We choose all instances of Science subset (Rafailov et al., 2024) and WizardLM Evol-Instruct (Xu et al., 2024) from the Tuluv2 mixture. Additionally we use all tasks from SuperNaturalInstructions (Wang et al., 2022) and sample a maximum of $100$ instances from each task. Lastly, we use all the instances from the CoT subset of the FLAN collection (Longpre et al., 2023b). Table 1 lists the datasets used in our experiments along with their statistics.
2.1 Pseudo-Code Generation
Super-Natural Instructions (SNI v2) dataset: (Wang et al., 2022): This is a collection of around $1600$ traditional NLP tasks formatted as input and output with prompts for instruction-tuning LLMs. We select $5$ pseudo-code demonstrations randomly from (Mishra et al., 2023)’s collection of $132$ tasks for the Generate step. As mentioned in Section 2 for the Evaluate and Repair steps, we prompt the Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2023) model with both pseudo-code instructions and natural languages instructions to generate the answers. We apply additional postprocessing such as splitting based on new line characters etc, to obtain the answer. We then use the exact-match metric to score outputs of the pseudo-code and natural language instructions and trigger the Repair step as needed.
We run the Repair twice with different temperature parameters and select the best program among the two runs. Table 9 in the Appendix reports the performance of the generated pseudo-code on the SNI v2 dataset after a Repair step where the ‘combined’ row represents the resulting performance. An example of a corrected pseudo-code program is presented in Appendix Figure 6.
GPT4-Alpaca, Code Alpaca WizardLM Evol-Instruct: For these datasets, we run the pseudo-code generation pipeline without Repair. These datasets usually have open-ended outputs and it is difficult to verify the correctness of the generated response unless we use a robust LLM-as-a-judge. Further, for these instances, forcing updates to the pseudo-code program using the Evaluate and Repair stages runs the risk of leaking output information into the pseudo-code program that can result in “CoT” like steps being included in the program. [^3] This isn’t a problem with SNI v2 because each task has many instances, and instance specific Repairs don’t generalize on other instances. While this isn’t detrimental and in fact likely to be helpful, we wanted to be careful and ensure we are able to study the impact of training with NL-equivalent pseudo-code instructions without any enrichment. Therefore, we only use the Generate part of our pipeline for these datasets. We select $5$ instances from each dataset and write the pseudo-code demonstrations ourselves. The instances are chosen to be diverse; for the the $1$-shot prompt, we randomly select one of these pseudo-code demonstrations.
Figure 2: Examples of different instructions types: (a) a distractor instruction from KCIF (Murthy et al., 2025) (b) nested instruction from ComplexBench (Wen et al., 2024) (c) simple sequential chain from ComplexBench (Wen et al., 2024) (d) two ‘levels’ of fine-grained constraints from FollowBench (Jiang et al., 2024b)
Filtering: We discard instances or tasks where the Mixtral model fails to generate valid pseudo-code instructions. This may occur if the pseudo-code is not enclosed within the ‘[PSEUDOCODE]’ and ‘[/PSEUDOCODE]’ tags, or if the generated pseudo-code is incomplete or malformed. Note that we also drop the corresponding instances from the natural language instruction set to ensure we have parallel training mixtures.
TuluV2 Science and CoT Data: We manually write pseudocode instructions for each of the $9$ datasets present in the CoT collection (Longpre et al., 2023b) and the Tulu V2 Science subsets (Ivison et al., 2023). The pseudo-code programs for each of these datasets are provided in the Appendix.
The data loop illustrated in Figure 1 follows an iterative generate–evaluate–repair paradigm for improving instruction-following behavior. Starting from a natural language (NL) instruction, the system generates a corresponding pseudo-code program using one-shot prompting. Both the NL instruction and the generated pseudo-code are independently evaluated via LLM inference and a scorer, yielding scores $S_{\text{NL}}$ and $S_{\text{PC}}$. When the pseudo-code underperforms (i.e., $S_{\text{NL}}$$S_{\text{PC}}$), targeted feedback is applied to repair and refine the program, thereby closing the loop. Only instruction-pseudo-code pairs that consistently satisfy the evaluation criteria are retained and stored along with their ground-truth outputs.
The effectiveness of this data loop is reflected in the evolution of dataset statistics of SNI v2 across iterations. The initial pass (Iteration 0) contains $1613$ instances generated with minimal filtering. After one full loop of evaluation and repair, Iteration 1 reduces the set to $662$ instances, indicating substantial pruning of low-quality or misaligned samples. A subsequent iteration further refines the dataset to $531$ instances. This steady reduction demonstrates that successive data loop iterations enforce increasingly strict quality thresholds, resulting in a high-precision dataset optimized for downstream training and evaluation.
Table 2: Comparison of Code-Prompted (CP) (Puerto et al., 2024), NL-tuned and pseudo-code (PC) tuned models on Instruction-Following, Mathematics, Commonsense & Reasoning tasks.
Instruction-Following Mathematics Common Sense & Reasoning Avg
Model IFEval KCIF
Comp.
Bench
Follow
Bench
AgentIF GSM8K
GSM8K
Plat.
Math
QA
ARC Hella. PiQA Wino.
Llama 3.1 8B-CP 0.32 0.09 0.25 0.20 0.37 0.14 0.14 0.18 0.90 0.24 0.76 0.44 0.34
Llama 3.1 8B-NL 0.39 0.08 0.42 0.40 0.32 0.33 0.33 0.21 0.82 0.71 0.77 0.66 0.45
Llama 3.1 8B-PC 0.46 0.13 0.46 0.47 0.42 0.61 0.63 0.29 0.89 0.71 0.82 0.61 0.54
Mistral 7B v0.3-CP 0.44 0.09 0.27 0.15 0.37 0.10 0.11 0.14 0.87 0.25 0.76 0.45 0.33
Mistral 7B v0.3-NL 0.35 0.08 0.35 0.36 0.32 0.12 0.13 0.21 0.74 0.53 0.65 0.40 0.35
Mistral 7B v0.3-PC 0.41 0.14 0.40 0.45 0.38 0.44 0.45 0.25 0.80 0.62 0.77 0.60 0.48
Qwen2.5 7B-CP 0.46 0.02 0.42 0.33 0.46 0.30 0.32 0.33 0.96 0.58 0.83 0.62 0.47
Qwen2.5 7B-NL 0.44 0.16 0.45 0.40 0.31 0.52 0.54 0.36 0.95 0.81 0.84 0.62 0.53
Qwen2.5 7B-PC 0.48 0.16 0.50 0.47 0.42 0.72 0.74 0.42 0.95 0.52 0.84 0.56 0.57
Granite 8B Code-CP 0.45 0.08 0.25 0.11 0.33 0.10 0.11 0.17 0.60 0.24 0.58 0.47 0.29
Granite 8B Code-NL 0.38 0.07 0.37 0.38 0.30 0.09 0.09 0.28 0.71 0.47 0.63 0.54 0.36
Granite 8B Code-PC 0.44 0.15 0.50 0.31 0.38 0.59 0.61 0.25 0.54 0.40 0.54 0.51 0.43
Granite 3.1 8B-CP 0.52 0.11 0.45 0.16 0.39 0.16 0.17 0.18 0.90 0.23 0.73 0.43 0.37
Granite 3.1 8B-NL 0.44 0.11 0.45 0.54 0.31 0.02 0.02 0.19 0.16 0.14 0.40 0.38 0.26
Granite 3.1 8B-PC 0.48 0.13 0.50 0.49 0.41 0.66 0.67 0.34 0.92 0.68 0.78 0.70 0.56
Olmo 3 7B-CP 0.46 0.09 0.40 0.40 0.42 0.25 0.25 0.23 0.91 0.07 0.72 0.49 0.39
Olmo 3 7B-NL 0.34 0.06 0.40 0.45 0.27 0.59 0.61 0.34 0.92 0.68 0.72 0.57 0.50
Olmo 3 7B-PC 0.36 0.05 0.48 0.52 0.44 0.73 0.75 0.37 0.92 0.66 0.74 0.55 0.55
3 Experiments
Through our experiments we seek to answer the following research questions: (i) Does instruction-tuning with pseudo-code help improve the performance of models on instruction-following tasks?, (ii) Does the introduction of pseudo-code during instruction-tuning affect the ability to acquire skills imparted during traditional instruction-tuning, (iii) How good is the quality of auto-generated pseudo code?, (iv) How does this method affect performance on code-tasks?
3.1 Models
Since we study the effect of instruction-tuning with and without pseudo-code instructions, we require unaligned base-models. After a careful review of available choices, we settled on the 787-8B parameter range models. Additionally, prior work by (Mishra et al., 2023) suggests that medium parameter-size models (eg: 6B-7B) were able to exploit pseudo-code better than smaller models. We experiment with the following $6$ base models which includes $1$ code base model: (i) meta-llama/Llama-3.1-8B, (ii) mistralai/Mistral-7B-v0.3, (iii) Qwen/Qwen2.5-7B, (iv) ibm-granite/granite-3.1-8b-base, (v) ibm-granite/granite-8b-code-base-128k, and (vi) allenai/Olmo-3-1025-7B.
Training: All models were SFT trained on a full NVIDIA A100 80GB node with Flash Attention $2$ (Dao, 2024) and bf16 precision using the Dolomite Engine Framework. [^4] https://github.com/IBM/dolomite-engine All models were trained for $2$ epochs using the PyTorch’s implementation of AdamW  (Loshchilov and Hutter, 2019) as the optimizer, with a starting learning rate of 1e-5 and linear decay. Models were trained with two parallel training mixtures - one with NL instructions and the other augmented with pseudo-code programs. We also compare the trained models against the inference-time method proposed by Puerto et al. (2024), which uses in-context examples to guide intermediate pseudo-code generation. In our evaluation of this method, we use hand-crafted, one-shot, task-specific demonstrations on base models to control for training data and keeping the comparisons meaningful.
3.2 Evaluation Tasks
We conduct experiments on four main types of tasks: (i) Instruction-Following, (ii) Common Sense and Reasoning, (iii) Mathematics and, (iv) Code Tasks.
Instruction Following: We conduct extensive experiments on instruction-following benchmarks, including IFEval (Zhou et al., 2023), KCIF (Murthy et al., 2025), ComplexBench (Wen et al., 2024), FollowBench (Jiang et al., 2024b), and a subset of AgentIF (Qi et al., 2025). [^5] Our training mixture does not include tool-calling and we remove those tasks from AgentIF.
IFEval and KCIF use non-LLM-based scoring functions to evaluate instruction-following performance. In contrast, ComplexBench, FollowBench, and AgentIF use powerful LLMs as judges. ComplexBench enables the study of different instruction compositions (e.g., conjunctions, chains, conditionals), while FollowBench reports performance aggregated by task categories (e.g., instructions affecting response styles, situations, or formats). AgentIF contains prompts from real-world tasks covering diverse constraint types. Notably, IFEval assesses only whether lexical and formatting instructions are followed, without considering task completion, whereas all other datasets condition instruction-following on core-task performance.
Common Sense and Reasoning Tasks: To evaluate whether our method affects skills acquired by vanilla natural-language SFT, we study performance on common benchmarks, including Reasoning tasks (HellaSwag (Zellers et al., 2019), PiQA (Bisk et al., 2020), Winogrande (Sakaguchi et al., 2021), and ARC (Clark et al., 2018)) and Mathematical tasks (GSM8K (Cobbe et al., 2021), GSM8K Platinum (Vendrow et al., 2025), and MathQA (Amini et al., 2019)).
Code Tasks: Since our SFT training includes Python-like pseudo-code, we investigate its impact on HumanEvalPack code benchmarks (Muennighoff et al., 2024). HumanEvalPack consists of three main categories—Code Repair, Code Explanation, and Code Synthesis—across six programming languages: Python, JavaScript, Java, Go, C++, and Rust. Although our training mixture included only a small proportion of code-related tasks (15% of the training data), most modern LLMs are pre-trained on substantial amounts of code.
Metrics: We post-process model outputs and use official scoring scripts where available. IFEval uses custom scorers for each instruction and aggregates results at both the prompt and instruction levels. ComplexBench, FollowBench, and AgentIF aggregate LLM-judge scores by task category. KCIF, as well as the Common Sense and Reasoning tasks, are measured using exact-match scores.
Table 3: ComplexBench: Performance of models on different instruction compositions (And, Chain, Selection) along with levels of nesting (11,22, $\geq 3$) where applicable.
And Chain Selection Selection and Chain
1 2 1 2 $\geq$3 2 $\geq$3
Llama 3.1 8B-NL 0.61 0.37 0.30 0.44 0.33 0.25 0.29 0.24
Llama 3.1 8B-PC 0.63 0.39 0.37 0.40 0.35 0.31 0.44 0.31
Mistral 7B v0.3-NL 0.51 0.29 0.25 0.35 0.25 0.23 0.25 0.20
Mistral 7B v0.3-PC 0.56 0.35 0.32 0.39 0.31 0.27 0.24 0.22
Qwen2.5 7B-NL 0.61 0.40 0.37 0.47 0.37 0.24 0.37 0.24
Qwen2.5 7B-PC 0.64 0.48 0.41 0.47 0.45 0.38 0.42 0.31
Granite 8B Code-NL 0.53 0.30 0.28 0.44 0.29 0.18 0.19 0.22
Granite 8B Code-PC 0.58 0.36 0.34 0.38 0.32 0.26 0.27 0.19
Granite 3.1 8B-NL 0.62 0.41 0.36 0.46 0.35 0.24 0.39 0.26
Granite 3.1 8B-PC 0.65 0.44 0.42 0.47 0.42 0.41 0.33 0.30
Olmo 3 7B-NL 0.56 0.33 0.33 0.35 0.32 0.21 0.28 0.24
Olmo 3 7B-NL-PC 0.64 0.45 0.41 0.45 0.37 0.30 0.37 0.32
3.2.1 Evaluation Setup and Metrics
Evaluation Setup: All tasks are evaluated under generative settings; this means that in the case of multiple choice questions models are required to verbalize the answer option  (Mishra et al., 2023; Li et al., 2024b). Further, all tasks are evaluated in zero-shot settings with greedy decoding; prompts for all mathematics, reasoning and code tasks encourage thinking step by step. Prompts for the instruction-following tasks are used as included in the original datasets. We use vLLM (Kwon et al., 2023) to speed up inference on the Mathematics, Reasoning and Code Tasks.
Output Post Processing: We develop robust task-specific output-processing pipelines for both NL and pseudo-code models. (See Appendix Section H for more details). All tasks are evaluated after output post-processing.
3.3 Results
Overall Results: As shown in Table 2, pseudo-code (PC) training improves instruction-following performance across all models, with PC variants outperforming natural-language (NL)-trained baselines on instruction-following tasks. The gains are also substantial on mathematical reasoning taskswhile for commonsense reasoning, results are more mixed: Qwen2.5 models do not exhibit a clear pattern, Granite 3.1 and Mistral 7B v0.3 show significant improvements with pseudo-code. We hypothesize that for instructions involving multi-step constraints, pseudo-code reduces ambiguity and clarifies task structure, which aligns with improved reasoning and instruction-following performance. In contrast, commonsense and reasoning tasks typically involve simpler instructions, which may explain the relatively smaller or mixed gains on these benchmarks. Notably, CP (code-prompted) baselines underperform on most tasks with the exception of ARC scores, suggesting training-time integration is more effective than inference-time prompting. We provide qualitative example outputs for all tasks in Appendix Section E.
3.3.1 Instruction-Following
We now present a deeper study on the following aspects of instruction-following: (i) effect of distractor instructions ,(ii) nested instructions, (iii) formatting instructions, (iv) fine-grained instruction-following. See Figure 2 for examples of these instruction types.
Instructions with Distractors: The KCIF benchmark  (Murthy et al., 2025) studies the performance of models on answer-modifying instructions when compared against regular question-answering (printing the correct-answer label/text). We use this dataset to study the performance of models on ‘distractors’ (instructions that should have no effect) as well as instructions that are conditional on a model’s ability to answer a QA task correctly. All models (with the exception of Qwen2.5 on distractors) universally perform better on both categories of instructions when trained with pseudo-code instructions (See Table 7 in appedidx).
Chained and Nested Instructions: We use ComplexBench  (Wen et al., 2024) to study how models perform in the presence of conjunctions (And), sequential instructions that are simple chains (Chain), conditional sequential operations (Sequential) at varying depths of nesting and, combinations of such instructions (eg. selections and chain). As can be seen in Table 3, models trained with pseudo-code instructions consistently perform better than those trained with natural language instructions. This holds true for almost all instruction-categories as well as nesting-depths. It is also interesting to note that doing well on a shallow nesting-depth does not automatically mean that the performance trends follow on deeper nesting-depths (eg: LLama 3.1 8B and Granite 8B Code NL models on Selection tasks at nesting levels $1$ vs the rest).
Table 4: Performance of models on formatting instructions from ComplexBench.
Model JSON Format
Markdown
Format
Bullet
Format
Punctuation Length Start with End with Template
Llama 3.1 8B-NL 0.13 0.66 0.45 0.43 0.27 0.17 0.34 0.39
Llama 3.1 8B-PC 0.18 0.81 0.57 0.49 0.26 0.18 0.27 0.40
Mistral 7B v0.3-NL 0.11 0.58 0.45 0.23 0.24 0.15 0.25 0.23
Mistral 7B v0.3-PC 0.16 0.68 0.52 0.29 0.27 0.19 0.27 0.34
Qwen2.5 7B-NL 0.18 0.66 0.49 0.38 0.30 0.15 0.25 0.48
Qwen2.5 7B-PC 0.18 0.74 0.58 0.40 0.28 0.19 0.28 0.47
Granite 8B Code-NL 0.13 0.68 0.42 0.35 0.28 0.14 0.30 0.28
Granite 8B Code-PC 0.18 0.70 0.53 0.42 0.29 0.21 0.31 0.33
Granite 3.1 8B-NL 0.18 0.64 0.53 0.32 0.29 0.18 0.34 0.40
Granite 3.1 8B-PC 0.22 0.77 0.59 0.46 0.27 0.18 0.28 0.48
Olmo 3 7B-NL 0.09 0.53 0.50 0.25 0.27 0.17 0.23 0.24
Olmo 3 7B-PC 0.13 0.70 0.58 0.40 0.29 0.15 0.33 0.41
Formatting Instructions: Using the data in ComplexBench, we study the performance of models on formatting instructions - as an be seen in Table 4 all models trained with pseudo-code instructions outperform their natural-language counterparts except on instructions that require length control. We think that since we require LLMs to also generate pseudo-code models before the final answer it might require specialized training to recalibrate length controls. Performance gains on JSON and Bullet formatting are especially notable ($11-27\%$ relative gains).
Table 5: Performance of models on fine-grained constraints from FollowBench
Hard Satisfaction Rate (HSR) % Soft Satisfaction Rate (SSR) %
Model L1 L2 L3 L4 L5 Avg. L1 L2 L3 L4 L5 Avg.
Llama 3.1 8B -NL 51.46 40.31 34.00 19.96 8.28 30.80 51.46 53.55 56.94 45.35 38.99 49.26
Llama 3.1 8B -PC 58.14 45.48 36.22 28.60 23.33 38.35 58.14 52.52 54.23 49.26 49.70 52.77
Mistral 7B v0.3 -NL 47.06 29.79 29.84 16.79 14.25 27.55 47.06 44.78 48.26 41.73 40.31 44.43
Mistral 7B v0.3 -PC 62.70 46.90 37.82 14.78 23.46 37.13 62.70 60.16 56.50 46.92 50.33 55.32
Qwen2.5 7B -NL 50.21 34.80 32.75 25.84 14.92 31.70 50.21 52.02 53.04 46.98 42.67 48.98
Qwen2.5 7B -PC 61.32 50.72 35.31 27.45 17.81 38.52 61.32 60.62 56.42 52.33 50.31 56.20
Granite 8B Code -NL 44.92 36.70 33.13 21.01 13.61 29.87 44.92 50.35 50.36 40.70 39.44 45.15
Granite 8B Code-PC 43.86 30.42 25.18 12.68 10.28 24.48 43.86 41.33 40.84 31.12 31.25 37.68
Granite 3.1 8B -NL 67.38 55.59 49.77 38.04 20.69 46.29 67.38 66.51 66.17 60.26 52.68 62.60
Granite 3.1 8B -PC 61.01 51.46 40.38 38.65 19.05 42.11 61.01 61.01 55.78 56.56 49.06 56.68
Olmo 7B -NL 54.16 51.68 33.25 24.21 17.86 36.23 54.16 62.02 54.64 50.18 45.62 53.32
Olmo 7B -PC 66.76 57.76 42.44 30.61 22.10 43.93 66.76 66.89 61.71 52.80 49.764 59.58
Fine-grained Instructions: FollowBench (Jiang et al., 2024b) includes instruction tasks where fine-grained constraints are incrementally introduced at each ‘level’. The dataset consists of $5$ levels of constraints applied to different task types - ‘style’, ‘situations’, ‘content’, ‘format’ and ‘examples’. Models are scored by assessing how well all constraints are satisfied for each instruction (hard-constraint satisfaction rate) and by assessing whether a particular constraint is satisfied, regardless of whether co-occurring constraints are satisfied are not (soft-constraint satisfaction rate).
We report the results of this experiment in Table 5. We find that with the exception of Granite 8B Code and Granite 3.1 8B models, [^6] We inspected the results from the Granite models but could not find a pattern that explains this behavior. all other models benefit from training with pseudo-code instructions. Further, performance gains when pseudo-code training is used are especially significant at higher (more complex) levels of constraints.
3.3.2 Quality of Pseudo-Code
Table 6: Performance of trained models when prefiled with human-generated pseudo-code vs auto-generated pseudo-code. Values represent accuracy or task-specific metrics.
Model Pseudocode ARC GSM8K GSM8K Plat. PIQA MathQA Hella. Wino. IFEval ComplexBench Avg
Qwen2.5 7B-PC Human 0.85 0.80 0.81 0.50 0.75 0.60 0.76 0.74 0.61 0.71
Model 0.94 0.65 0.85 0.60 0.35 0.75 0.55 0.64 0.62 0.66
Mistral-7B-v0.3-PC Human 0.80 0.45 0.25 0.80 0.15 0.55 0.55 0.51 0.45 0.50
Model 0.54 0.30 0.65 0.60 0.05 0.55 0.60 0.46 0.47 0.47
Granite 8B Code-PC Human 0.75 0.60 0.45 0.65 0.05 0.10 0.55 0.46 0.49 0.46
Model 0.07 0.14 0.65 0.63 0.05 0.10 0.55 0.31 0.52 0.34
Granite 3.1 8B-PC Human 0.95 0.55 0.75 0.85 0.20 0.75 0.85 0.69 0.67 0.70
Model 0.57 0.45 0.70 0.60 0.25 0.75 0.85 0.59 0.65 0.60
Llama 3.1-8B-PC Human 0.95 0.65 0.15 0.85 0.15 0.71 0.70 0.58 0.64 0.60
Model 0.89 0.45 0.40 0.60 0.30 0.70 0.50 0.52 0.44 0.53
Olmo 3 7B-PC Human 0.55 0.70 0.30 0.75 0.20 0.75 0.80 0.58 0.66 0.59
Model 0.25 0.60 0.60 0.70 0.25 0.75 0.75 0.56 0.56 0.56
To assess the quality of our automatically generated pseudo-code, we compared model-generated and human-written examples by using them as prefill prompts for trained models, across 20 instances per task totaling $200$ instances (Table 6).
Results vary by model and task: for example, on GSM8K Platinum, model-generated pseudo-code sometimes matches or outperforms human-written examples (e.g., 0.85 vs. 0.81 for Qwen2.5-7B), while on other tasks like Hellaswag, differences are minimal. Overall, these findings suggest that our automatic pseudo-code generation produces quality sufficient for the training gains observed in Table 2, though the gap between human and model pseudo-code on complex reasoning tasks (e.g., GSM8K Platinum for some models) indicates potential for further improvement in the generation pipeline.
3.3.3 Code Generation Tasks
Recall that during training we constrained pseudo-code representations to focus solely on task instructions, in order to prevent solution leakage in the pseudo-code itself (see Appendix Figure 20). Under this constraint, it is not obvious that pseudo-code instructions that do not include solution steps should benefit code-generation tasks. As with our decision to exclude CoT-style reasoning in other settings, this design choice isolates the effect of pseudo-code–based instruction training on model performance.
We evaluate models on four HumanEvalPack tasks that probe different aspects of code-generation reasoning. FixTests requires repairing incorrect code using only failing unit tests. FixDocs removes test supervision, challenging models to fix code using natural language documentation alone. Synthesize is a standard code-generation task in which programs are produced from problem descriptions. ExplainSynthesize evaluates whether models can generate correct implementations from intermediate explanations, testing more abstract reasoning.
Somewhat surprisingly, Table 8 (Appendix) shows that, on average, training with pseudo-code instructions improves performance over natural-language prompting across HumanEvalPack tasks. Granite 8B Code exhibits the largest average gain, increasing from 0.301 (NL) to 0.365 (pseudo-code), with especially large improvements on FixDocs and Synthesize in Rust. While gains vary by model and task, these results suggest that re-expressing instructions in a pseudo-code form can positively affect code-generation performance, even when the representation is limited to instruction specification rather than solution guidance.
4 Related Work
Our work is most closely related to the creation of intermediate thinking/reasoning representations LLMs. Most often, these are applied at inference time - styles include simple prompts to elicit thinking  (Kojima et al., 2022), instructions that guide LLMs to breakdown tasks and plan  (Xu et al., 2023), structured representation and meta chain-of-thought (eg: Tree Of Thoughts  (Yao et al., 2023a), Graph of Thoughts  (Besta et al., 2024), Table of Thoughts (Sun et al., 2025), Buffer of Thoughts (Yang et al., 2024b)), specialized representations for tasks (eg: Tabular Reasoning  (Wang et al., 2024), Numerical Reasoning  (Chen et al., 2023), Code Generation  (Li et al., 2025), Conditional QA  (Lin et al., 2024)), etc. Additionally, prompting strategies that use code-like representation to aid thinking have also been developed  (Mishra et al., 2023; Li et al., 2024a; Puerto et al., 2024). Our work extends this body of work on the use of code-like representations - specifically, instead of relying on tedious few-shot demonstrations to elicit code-generation during inference, we fine-tune models to do this as part of a ‘thinking’ step before returning the answer. We mirror the approach taken by  (Mishra et al., 2023) and  (Puerto et al., 2024) of using pseudo-code as ‘instructions’ during inference and do not rely on the use of a code-emulator for line-by-line attempted execution used in Chain-of-Code (Li et al., 2024a). Thus, our models do not require external code-execution environments or emulators can be used as drop-in replacements for their corresponding NL-instruction-tuned model counterparts.
5 Conclusion and Future Work
In this paper, we explored an instruction-tuning strategy in which models are trained to first re-express a given instruction in pseudo-code before attempting to solve the task. We modified a subset of the publicly available TuluV2 dataset (Ivison et al., 2023) and fine-tuned six publicly available base models. Extensive experiments across a range of instruction-following benchmarks demonstrate that fine-tuning with pseudo-code generally outperforms training with natural language (NL) instructions alone.
Our findings suggest several promising directions for future work. To isolate the effect of pseudo-code, we deliberately ensured that the pseudo-code representations did not include chain-of-thought or other instruction-enrichment strategies (Qi et al., 2024; Lou et al., 2024). However, Table 2 shows that for certain model-task combinations, particularly in commonsense reasoning, training with NL instructions can sometimes be more effective. This points to potential hybrid approaches, such as combining NL and pseudo-code instructions, exploring mixture-of-experts architectures (Jiang et al., 2024a), applying model merging techniques (Yang et al., 2024a), or extending pseudo-code guidance to pretraining.
Further, training models to think and plan in code-like representations may enable more powerful applications, especially when combined with frameworks like LMQL (Beurer-Kellner et al., 2023), which support both LLM-executable and code-executable programs with control flow.
6 Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Overview of Appendix
Our appendix is organized as follows. We begin by including additional results studying the performance of models in the presence of distractors (Appendix Section B). We then include additional results studying the performance of our method on code tasks in Appendix Section C, followed by PC and NL training samples in Appendix Section D. We include sample outputs from trained models in Appendix Section E. We then include the quantitative and qualitative results of running the Repair stage in Appendix Section F. We include data specific instances pseudo-code programs used in the creation of training data in Appendix Section G and then provide details on how model outputs are processed for evaluation in Appendix Section H. Finally, we include some details on statistical significance (Appendix Section I) and conclude the Appendix by discussing limitations of our work in Appendix Section J.
Table 7: KCIF: Performance comparison on non-distractor (EM) and distractor instructions (EM-D). On the right hand side of ‘||’ separator, Performance drop as compared to printing the correct answer label (PCA Label) or text (PCA) on answer-modifying instructions (String/Numeric/Format manipulation, Operations on List (OOL) – conditioned on the answering the question correctly as well as independent operations on list). Barring the PCA Label column, for all columns to its right - lower scores are better.
Instruction Category
Model
EM
EM-D
PCA label PCA String Numeric Format
OOL
Cond.
OOL Avg
LLama 3.1 8B-NL 0.04 0.09 0.34 -0.29 -0.34 -0.34 -0.34 -0.28 -0.34 -0.33 / -96.2%
LLama 3.1 8B-PC 0.09 0.20 0.38 -0.19 -0.36 -0.38 -0.25 -0.25 -0.38 -0.32 / -85.3%
Mistral 7B v0.3-NL 0.04 0.15 0.32 -0.16 -0.3 -0.31 -0.32 -0.26 -0.32 -0.3 / -95.0%
Mistral-7B v0.3-PC 0.08 0.27 0.37 -0.15 -0.37 -0.37 -0.31 -0.23 -0.37 -0.33 / -88.4%
Qwen2.5 7B-NL 0.09 0.30 0.41 -0.12 -0.4 -0.4 -0.31 -0.28 -0.41 -0.36 / -87.5%
Qwen2.5 7B-PC 0.11 0.26 0.46 -0.3 -0.44 -0.46 -0.31 -0.29 -0.44 -0.39 / -83.7%
Granite 8B Code-NL 0.04 0.12 0.29 -0.19 -0.28 -0.29 -0.27 -0.29 -0.29 -0.28 / -97.2%
Granite 8B Code-PC 0.10 0.26 0.33 -0.14 -0.31 -0.33 -0.19 -0.19 -0.32 -0.27 / -81.3%
Granite 3.1 8B-NL 0.07 0.18 0.39 -0.29 -0.38 -0.39 -0.37 -0.31 -0.39 -0.37 / -94.6%
Granite 3.1 8B-PC 0.07 0.25 0.43 -0.27 -0.43 -0.43 -0.39 -0.32 -0.43 -0.4 / -92.9%
Olmo 3 7B-NL 0.04 0.10 0.42 -0.39 -0.42 -0.42 -0.41 -0.36 -0.42 -0.41 / -96.7%
Olmo 3 7B-PC 0.04 0.20 0.46 -0.43 -0.46 -0.46 -0.46 -0.38 -0.46 -0.44 / -96.3%
Appendix B Performance on Instructions with Distractors
The KCIF benchmark  (Murthy et al., 2025) studies the performance of models on answer-modifying instructions when compared against regular question-answering (printing the correct-answer label/text). We use this dataset to study the performance of models on ‘distractors’ (instructions that should have no effect) as well as instructions that are conditional on a model’s ability to answer a QA task correctly. All models (with the exception of Qwen2.5 on distractors) universally perform better on both categories of instructions when trained with pseudo-code instructions.
Instructions with Distractors: Table 7 presents two sets of results: On the left hand side of ‘||’ separator, the table compares the exact-match performance of models (after output post-processing) on instructions that do not have distractor instructions [^7] Eg.: an instruction that asks to increment any numerical values when the answers are non-numeric text. (EM) vs. ones that do (EM-D). As can be seen, all models (with the exception of Qwen2.5 on distractors) universally perform better on both categories of instructions when trained with pseudo-code instructions.
Answer-Modifying Instructions: On the right hand side of Table 7, we compare how instruction-compositions applied on the answering task affect performance. Specifically, given a traditional multi-choice knowledge or reasoning task, models first need to generate the correct answer and then apply the instruction that suitable modifies that answer.
As compared to the baseline task of question-answering by printing the correct answer label (PCA Label), asking models to print the correct answer text (PCA) or applying answer-modifying instructions (string, numeric or format manipulations, operations on lists regardless of whether they are conditional on the answering task) result in a steep drop in performance. It is noteworthy, that while all models (independent of the training paradigm) have steep deterioration in performance, training with pseudo-code instructions results in a lesser drop in performance.
Appendix C Performance on Code Tasks
Table 8: Model Performance Scores Across All Programming Languages (NL and PC-COT) [^8] The results reported in an earlier version of this paper were affected by a bug in the evaluation code. The issue has been identified and corrected, and all results have been recomputed accordingly.
FixTests FixDocs Synthesize ExplainSynth
Model JS Py C++ Rust JS Py C++ Rust JS Py C++ Rust JS Py C++ Rust Avg
Llama 3.1 8B-NL 0.213 0.220 0.116 0.140 0.390 0.445 0.262 0.165 0.341 0.372 0.268 0.189 0.049 0.067 0.037 0.049 0.208
Llama 3.1 8B-PC 0.104 0.226 0.146 0.006 0.317 0.427 0.280 0.159 0.378 0.445 0.299 0.213 0.079 0.049 0.037 0.037 0.200
Mistral 7B v0.3-NL 0.012 0.085 0.067 0.110 0.043 0.159 0.055 0.226 0.134 0.232 0.213 0.085 0.055 0.043 0.061 0.018 0.097
Mistral 7B v0.3-PC 0.043 0.079 0.049 0.043 0.189 0.165 0.146 0.067 0.213 0.250 0.195 0.067 0.067 0.073 0.067 0.006 0.118
Qwen2.5 7B-NL 0.476 0.500 0.329 0.146 0.622 0.683 0.616 0.183 0.543 0.610 0.488 0.183 0.159 0.140 0.085 0.183 0.372
Qwen2.5 7B-PC 0.134 0.500 0.335 0.195 0.677 0.683 0.573 0.476 0.628 0.652 0.500 0.494 0.183 0.171 0.073 0.146 0.401
Granite 8B Code-NL 0.232 0.378 0.274 0.226 0.518 0.622 0.524 0.201 0.476 0.476 0.415 0.159 0.079 0.098 0.079 0.055 0.301
Granite 8B Code-PC 0.323 0.409 0.274 0.262 0.616 0.659 0.573 0.463 0.561 0.555 0.427 0.402 0.091 0.073 0.085 0.061 0.365
Granite 3.1 8B-NL 0.268 0.256 0.195 0.226 0.372 0.470 0.390 0.250 0.415 0.463 0.390 0.122 0.079 0.067 0.024 0.030 0.251
Granite 3.1 8B-PC 0.195 0.293 0.189 0.165 0.390 0.457 0.372 0.244 0.366 0.415 0.390 0.207 0.030 0.043 0.030 0.030 0.239
Since our SFT training includes Python-like pseudo-code, we investigate its impact on HumanEvalPack code benchmarks (Muennighoff et al., 2024). HumanEvalPack consists of three main categories—Code Repair, Code Explanation, and Code Synthesis—across six programming languages: Python, JavaScript, Java, Go, C++, and Rust. Although our training mixture included only a small proportion of code-related tasks (15% of the training data), most modern LLMs are pre-trained on substantial amounts of code.
We evaluate models on four HumanEvalPack tasks that test code-generation reasoning and robustness. FixTests requires repairing incorrect code using only failing unit tests. FixDocs removes test supervision, challenging the model to fix code from natural language documentation. Synthesize is the standard code-generation task, producing programs from problem descriptions. ExplainSynthesize evaluates whether models can generate correct implementations from intermediate explanations, testing abstract reasoning.
Table 8 shows that, on average, training with pseudo-code instructions improves performance over natural-language prompting across HumanEvalPack tasks. Granite 8B Code exhibits the largest average gain, increasing from 0.301 (NL) to 0.365 (pseudo-code), with especially large improvements on FixDocs and Synthesize in Rust. While gains vary by model and task, these results suggest that re-expressing instructions in a pseudo-code form can positively affect code-generation performance, even when the representation is limited to instruction specification rather than solution guidance.
Appendix D Training Sample
Figure 3 and 4 shows training samples for PC and NL respectively.
Figure 3: PC Training Sample
Write a short story in third person narration about a protagonist who has to make an important career decision. Generate pseudocode chain-of-thought reasoning first and then generate the solution. Input
[PSEUDOCODE]
def code\_write\_short\_story\_career\_decision(input: str) -> str:
    Write a short story in third person narration about a protagonist who has to make an important career decision.
    Parameters:
        str: input
    Returns:
        str: code
    # this function takes the input string as an input
    protagonist\_name = extract\_name(input)
    # this function generates a code to write a short story in third person narration
    # about a protagonist who has to make an important career decision
    code = generate\_code(input={\"protagonist\_name\": protagonist\_name, \"task\": \"write short story about
    career decision\"}, story\_type=\"third person narration\")
    return code
>>> code\_write\_short\_story\_career\_decision(input)
[/PSEUDOCODE]
Sophie sat at her desk, staring blankly at the computer screen. Her mind was racing as she weighed the options in front of
her. She had been offered a promotion to a senior management position within her company, but it meant moving across
the country and leaving her friends and family behind. At the same time, her dream job at a rival company had finally
opened up, and she had been invited for an interview.
The thought of leaving her comfortable life behind, to take a chance on something new and uncertain, made Sophie feel
anxious, but also excited. She had always been one to take risks, and it had paid off in the past. But now, with such a big
decision to make, Sophie couldn’t help but feel overwhelmed.
After much deliberation, Sophie made her choice. She packed her bags, said goodbye to her loved ones, and moved across the
country for the new job. It was a difficult first few months, adjusting to a new city and a new company culture. But as time
went by, Sophie realized that she had made the best decision for her career. The job was challenging, but also rewarding,
and she was surrounded by passionate, driven colleagues who inspired her to grow and develop.
In the end, Sophie’s bold move paid off. She was thriving in her new role, and had made new friends and connections in her
new city. Looking back, Sophie was grateful for the opportunity to take a risk and make a change, and was confident that she
had made the right decision for her career.
Output
Figure 4: NL Training Sample
Write a short story in third person narration about a protagonist who has to make an important career decision. Input
Sophie sat at her desk, staring blankly at the computer screen. Her mind was racing as she weighed the options in front of
her. She had been offered a promotion to a senior management position within her company, but it meant moving across
the country and leaving her friends and family behind. At the same time, her dream job at a rival company had finally
opened up, and she had been invited for an interview.
The thought of leaving her comfortable life behind, to take a chance on something new and uncertain, made Sophie feel
anxious, but also excited. She had always been one to take risks, and it had paid off in the past. But now, with such a big
decision to make, Sophie couldn’t help but feel overwhelmed.
After much deliberation, Sophie made her choice. She packed her bags, said goodbye to her loved ones, and moved across the
country for the new job. It was a difficult first few months, adjusting to a new city and a new company culture. But as time
went by, Sophie realized that she had made the best decision for her career. The job was challenging, but also rewarding,
and she was surrounded by passionate, driven colleagues who inspired her to grow and develop.
In the end, Sophie’s bold move paid off. She was thriving in her new role, and had made new friends and connections in her
new city. Looking back, Sophie was grateful for the opportunity to take a risk and make a change, and was confident that she
had made the right decision for her career.
Output
Appendix E Qualitative Samples
E.1 IFEval
E.1.1 Example 1
Input Instruction:
Write a 300+ word summary of the wikipedia page "https://en.wikipedia.org/wiki/Raymond_III,_Count_of_Tripoli". Do not use any commas and highlight at least 3 sections that has titles in markdown format, for example *highlighted section part 1*, *highlighted section part 2*, *highlighted section part 3*.
Output from LLama 3.1 8B-PC
[PSEUDOCODE]
def generate_summary(url: str) -> str:
Generate a 300+ word summary of the wikipedia page.
Parameters:
str: url
Returns:
str: summary
# this function takes the url as input and generates a summary
summary = generate_wikipedia_summary(url=url, word_count=300)
# this function highlights the sections with titles in markdown format
summary = highlight_sections(summary)
return summary
>>> generate_summary(url)
[/PSEUDOCODE]
Raymond III, Count of Tripoli (c. 1140 - 1187), also known as Raymond of Antioch, was a prince of the County of Tripoli and the Kingdom of Jerusalem. He was the son of Count Pons of Tripoli and his wife, Melisende of Jerusalem. Raymond was a member of the House of Tripoli, a branch of the House of Anjou.
Raymond III is best known for his role in the Third Crusade, where he was a vassal of King Richard the Lionheart of England. He was one of the leaders of the Crusader army that defeated the Ayyubid forces at the Battle of Arsuf in 1191. He also participated in the Siege of Acre, which lasted for two years and is considered one of the bloodiest battles of the Crusades.
In addition to his military exploits, Raymond was also involved in diplomacy. He was one of the signatories of the Treaty of Ramla, which was signed in 1157 between the Kingdom of Jerusalem and the Fatimid Caliphate. He also served as a mediator between the Crusader states and the Byzantine Empire.
Raymond III was married twice. His first wife was Eschiva of Armenia, with whom he had two sons, Bohemond and Baldwin. After her death, he married Alice of Champagne, with whom he had two daughters, Melisende and Philippa.
Raymond III died in 1187 during the Battle of Cresson. His death was a major blow to the Crusader states, as he was one of their most capable leaders. He was succeeded by his son Bohemond, who was only a child at the time, and the County of Tripoli was ruled by regents until Bohemond came of age.
Output from LLama 3.1 8B-NL
Raymond III, Count of Tripoli (c. 1135 - 1187) was the ruler of the County of Tripoli from 1182 to 1187. He was the son of Raymond II of Tripoli and his wife Hodierna of Toron. Raymond III was involved in the Crusades and was captured by Saladin in 1187 during the Battle of Hattin. He was released after the payment of a ransom. Raymond III was also involved in the negotiations for the Third Crusade. Raymond III was married to Eschiva of Bures, with whom he had two children, Raymond and Alice.
Raymond III was involved in the Crusades and was captured by Saladin in 1187 during the Battle of Hattin. He was released after the payment of a ransom. Raymond III was also involved in the negotiations for the Third Crusade.
Raymond III, Count of Tripoli was the son of Raymond II of Tripoli and his wife Hodierna of Toron. Raymond III was involved in the Crusades and was captured by Saladin in 1187 during the Battle of Hattin. He was released after the payment of a ransom. Raymond III was also involved in the negotiations for the Third Crusade. Raymond III was married to Eschiva of Bures, with whom he had two children, Raymond and Alice.
Comment: As can be seen, Llama 3.1 8B-PC model first generated pseudo-code and then the final response. Not all aspects of the instruction are included in the generated pseudo-code in this example. Further, both models do not get the requirement of using section highlights correctly.
E.1.2 Example 2
Input Instruction:
Write a detailed review of the movie "The Social Network". Your entire response should be in English and all lower case (no capital letters whatsoever).
Output from Qwen2.5 7B-PC
def review_movie_social_network() -> str:
Write a detailed review of the movie "The Social Network". The entire response should be in English and all lower case (no capital letters whatsoever).
Returns:
str: response
# this function generates a detailed review of the movie "The Social Network"
review = generate_movie_review(movie_name="The Social Network", language="English", case="lower")
return review
>>> review_movie_social_network()
[/PSEUDOCODE]
the social network is a captivating and engaging film that tells the story of mark zuckerberg and the creation of facebook. the film is well-crafted and the acting is top-notch, with each character portrayed in a nuanced and realistic manner. the story is engaging and the pacing is well-balanced, making for a smooth and enjoyable viewing experience. the film also does a great job of exploring the ethical and social implications of facebook, making it a thought-provoking and relevant addition to the film canon. overall, the social network is a must-see for anyone interested in technology, entrepreneurship, or simply a well-told story.
Output from Qwen2.5 7B-PC
The Social Network is a 2010 biographical drama film directed by David Fincher, based on the book "The Accidental Billionaires: The Founding of Facebook, the Early Years of Mark Zuckerberg" by Ben Mezrich. The film stars Jesse Eisenberg as Mark Zuckerberg, Andrew Garfield as Eduardo Saverin, and Justin Timberlake as Sean Parker. The movie tells the story of the creation of Facebook and the legal battle that ensued between Zuckerberg and his former business partners.
The film is well-crafted and engaging, with a strong performance from Jesse Eisenberg as Zuckerberg. Eisenberg perfectly captures the character’s mix of intelligence, charm, and arrogance, making him both likable and relatable. The supporting cast is also strong, with Andrew Garfield delivering a nuanced performance as Eduardo Saverin, and Justin Timberlake as the charismatic Sean Parker.
The Social Network is a well-paced film that keeps the audience engaged throughout. The story is told in a non-linear fashion, jumping back and forth between different time periods, which adds to the tension and keeps the audience guessing. The film also does a good job of balancing the drama and the humor, making it a well-rounded and enjoyable experience.
One of the highlights of the film is the portrayal of the legal battle between Zuckerberg and his former business partners. The tension and drama of the courtroom scenes are palpable, and the film does a good job of showing the impact that the lawsuit had on Zuckerberg’s life and career.
Overall, The Social Network is a well-crafted and engaging film that tells a compelling story. The performances are strong, the pacing is good, and the film does a good job of balancing the drama and the humor. Highly recommended.
Comment: As can be seen, Qwen2.5 7B-PC model first generated pseudo-code and then the final response. It follows the instruction correctly, while in this particular example Qwen2.5 7B-NL fails to respond in lowercase.
E.2 ARC
E.2.1 Example 1
Input :
When the soil is saturated, earthworms move to the top of the soil to obtain oxygen. This behavior is evidence of which biological concept in earthworms?
Options:
A. learned behavior
B. migration instinct
C. response to stimuli
D. reproductive strategy
Ground Truth:
Ground Truth:
C
Processed Ground Truth:
C
Output from Granite 3.1 8B-NL :
Generated Output:
C. Response:C.
Processed Generated Output:
C
E.2.2 Example 2
Input :
Which statement best explains why photosynthesis is the foundation of most food webs?
Options:
A. Sunlight is the source of energy for nearly all ecosystems.
B. Most ecosystems are found on land instead of in water.
C. Carbon dioxide is more available than other gases.
D. The producers in all ecosystems are plants.
Ground Truth:
Ground Truth:
A
Processed Ground Truth:
A
Output from Granite 3.1 8B-PC :
Generated Output:
[PSEUDOCODE]
def generate_answer(question: str, options: List[str]) -> str:
Given a question and 4 options, generate one of A, B, C, or D which best answers the question.
Parameters:
str: question
List[str]: options
Returns:
str: answer
# read the question and the options
# use our knowledge to select the best option
# return the selected option
>>> generate_answer(Given a question and 4 options, generate one of A, B, C, or D which best answers the question.
Question: Which statement best explains why photosynthesis is the foundation of most food webs?
Options:
A. Sunlight is the source of energy for nearly all ecosystems.
B. Most ecosystems are found on land instead of in water.
C. Carbon dioxide is more available than other gases.
D. The producers in all ecosystems are plants.
Answer: A.)
[/[PSEUDOCODE]]
Sunlight is the source of energy for nearly all ecosystems."
Processed Generated Output:
A
E.3 GSM8K
E.3.1 Example 1
Input :
Toulouse has twice as many sheep as Charleston. Charleston has 4 times as many sheep as Seattle. How many sheep do Toulouse, Charleston, and Seattle have together if Seattle has 20 sheep?
Ground Truth:
Ground Truth:
[…]
Processed Ground Truth:
260
Output from Llama 3.1 8B-NL :
Generated Output:
C. Response:4
Processed Generated Output:
4
E.3.2 Example 2
Input :
Eliza’s rate per hour for the first 40 hours she works each week is \$10. She also receives an overtime pay of 1.2 times her regular hourly rate. If Eliza worked for 45 hours this week, how much are her earnings for this week?’
Ground Truth:
Ground Truth:
[…]
Processed Ground Truth:
460’
Output from Llama 3.1 8B-PC :
Generated Output:
[PSEUDOCODE]
def eliza_weekly_earnings() -> str:
Eliza’s rate per hour for the first 40 hours she works each week is \$10. She also receives an overtime pay of 1.2 times her regular hourly rate. If Eliza worked for 45 hours this week, how much are her earnings for this week?
Returns:
str: response
# this function takes in the number of hours Eliza worked this week
hours_worked = 45
# this function calculates the regular pay for the first 40 hours
regular_pay = calculate_regular_pay(hours=40, hourly_rate=10)
# this function calculates the overtime pay
overtime_pay = calculate_overtime_pay(hours=hours_worked - 40, hourly_rate=10)
# this function calculates the total pay
total_pay = regular_pay + overtime_pay
return \"Eliza’s earnings for this week is \$\" + str(total_pay)
>>> eliza_weekly_earnings()
[/PSEUDOCODE]
For the first 40 hours, Eliza earns \$10 x 40 = \$400.
For the overtime hours, Eliza earns 1.2 x \$10 = \$12 per hour.
Therefore, for the 5 overtime hours, Eliza earns \$12 x 5 = \$60.
Thus, Eliza’s total earnings for the week is \$400 + \$60 = \$460."
Processed Generated Output:
460
E.4 GSM8K Platinum
E.4.1 Example 1
Input :
[…]
Ground Truth:
Ground Truth:
[…]
Processed Ground Truth:
7
Output from Granite 3.1 8B-NL :
Generated Output:
Response: 16
Processed Generated Output:
16
E.4.2 Example 2
Input :
Greg has an alarm set to ring three times a day as a reminder. When the alarm goes off, it continues to ring until Greg turns it off. The first time it went off today, it rang four times. The second time it went off, it rang for three times as long as the first time. The third time, it rang for half as long as the second time. How many times did the alarm ring in all?
Ground Truth:
Ground Truth:
[…]
Processed Ground Truth:
22
Output from Qwen2.5 7B-PC :
Generated Output:
The second time the alarm went off, it rang for 4 * 3 = 12 times. The third time the alarm went off, it rang for 12 / 2 = 6 times. In total, the alarm rang 4 + 12 + 6 = 22 times.
[PSEUDOCODE]
def answer_question(question: str) -> str:
This function answers the given question
Parameters:
str: question
Returns:
str: answer
# this function extracts the question from the input
question = extract_question(question)
# this function generates the answer to the question
answer = generate_answer(question)
return answer
>>> answer_question(Answer the following question. Generate pseudocode chain-of-thought reasoning first and then generate the solution.
Question: Greg has an alarm set to ring three times a day as a reminder. When the alarm goes off, it continues to ring until Greg turns it off. The first time it went off today, it rang four times. The second time it went off, it rang for three times as long as the first time. The third time, it rang for half as long as the second time. How many times did the alarm ring in all?
Answer: The second time the alarm went off, it rang for 4 * 3 = 12 times. The third time the alarm went off, it rang for 12 / 2 = 6 times. In total, the alarm rang 4 + 12 + 6 = 22 times.)
[/PSEUDOCODE]
Explanation: The first time it went off, it rang 4 times. The second time it went off, it rang for three times as long as the first time, or 4 * 3 = 12 times. The third time, it rang for half as long as the second time, or 12 / 2 = 6 times. In total, the alarm rang 4 + 12 + 6 = 22 times."
Processed Generated Output:
22
E.5 Hellaswag
E.5.1 Example 1
Input :
Sentence: Having an ice cream: A child sitting at restaurant table places his hands in a dessert. The child
Options:
0. eats the dessert and by the camera stares off into the distance.
1. places the hands back into the dessert.
2. uses his hands to eat the food.
3. stands and walks towards the camera.
Ground Truth:
Ground Truth:
2
Processed Ground Truth:
2
Output from Mistral 7B v0.3-NL :
Generated Output:
Response: 2
Processed Generated Output:
2
E.5.2 Example 2
Input :
Sentence: Cutting the grass: A man walks outside plugs his lawn mower in and gets ready to mow. He starts mowing his lawn casually going over all the grass. His dog
Options:
0. comes outside and starts to follow him a little while he’s closer to the house.
1. comes outside and takes a closer look at the man mowing, after he picks up the mower.
2. runs loose and feeds on his movements.
3. is having a seizure and starts to hop all over the lawn as he walks around.’
Ground Truth:
Ground Truth:
0
Processed Ground Truth:
0
Output from Mistral 7B v0.3-PC :
Generated Output:
0
[PSEUDOCODE]
def generate_correct_ending(sentence: str, options: List[str]) -> int:
Given a partial sentence and four options 0, 1, 2, and 3. Generate the option number corresponding to the correct ending.
Parameters:
sentence (str): sentence
options (List[str]): options
Returns:
int: option number
# this function checks if the partial sentence is complete or not
if is_sentence_complete(sentence):
# if the partial sentence is complete, then we return the option number corresponding to the correct ending
return get_correct_ending(sentence, options)
else:
# if the partial sentence is not complete, then we return -1
return -1
>>> generate_correct_ending(\"Cutting the grass: A man walks outside plugs his lawn mower in and gets ready to mow. He starts mowing his lawn casually going over all the grass. His dog
Options:
0. comes outside and starts to follow him a little while he’s closer to the house.
1. comes outside and takes a closer look at the man mowing, after he picks up the mower.
2. runs loose and feeds on his movements.
3. is having a seizure and starts to hop all over the lawn as he walks around.
Answer: 0\", [’0. comes outside a’, ’d starts to follow him a little while he\\’s closer to the house.’])
[/PSEUDOCODE]
0"
Processed Generated Output:
0
E.6 MathQA
E.6.1 Example 1
Input :
a can do a piece of work in 4 hours ; b and c together can do it in 3 hours , while a and c together can do it 2 hours . how long will b alone take to do it ?
Options:
a. 8 hours
b. 10 hours
c. 12 hours
d. 24 hours
e. none of these
Ground Truth:
Ground Truth:
c
Processed Ground Truth:
12hours
Output from Granite 8B Code-NL :
Generated Output:
cResponse: c
Processed Generated Output:
12hours
E.6.2 Example 2
Input :
the perimeter of a rhombus is 68 cm and one of its diagonals is 16 cm . find its area ?
Options:
a. 290
b. 240
c. 2870
d. 289
e. 292
Ground Truth:
Ground Truth:
b
Processed Ground Truth:
240
Output from Llama 3.1 8B-PC :
Generated Output:
b
Processed Generated Output:
240
E.7 PiQA
E.7.1 Example 1
Input :
ice box
Options:
0. will turn into a cooler if you add water to it
1. will turn into a cooler if you add soda to it
Ground Truth:
Ground Truth:
0
Processed Ground Truth:
0
Output from Granite 8B Code-NL :
Generated Output:
0
Processed Generated Output:
0
E.7.2 Example 2
Input :
To obtain meximum deliciousness of the nutella cookies
Options:
0. Serve with a glass of milk
1. Serve with a glass of water
Ground Truth:
Ground Truth:
0
Processed Ground Truth:
0
Output from Qwen2.5 7B-PC :
Generated Output:
0
[PSEUDOCODE]
def answer_question(question: str, option_0: str, option_1: str) -> int:
Given a question with two possible options 0 and 1. Generate either 0 or 1 which answers the question.
Parameters:
question (str): question
option_0 (str): option 0
option_1 (str): option 1
Returns:
int: answer
# this function checks if the question is a yes/no question
if is_yes_no_question(question):
# this function returns 0 if the answer is no
if answer_is_no(question):
return 0
# this function returns 1 if the answer is yes
elif answer_is_yes(question):
return 1
# this function checks if the question is a factual question
elif is_factual_question(question):
# this function returns 0 if the answer is false
if answer_is_false(question):
return 0
# this function returns 1 if the answer is true
elif answer_is_true(question):
return 1
else:
# this function returns 0 if the question cannot be answered
if question_cannot_be_answered(question):
return 0
# this function returns 1 if the question can be answered
elif question_can_be_answered(question):
return 1
>>> answer_question(\"To obtain meximum deliciousness of the nutella cookies
Options:
0. Serve with a glass of milk
1. Serve with a glass of water
Answer: 0\", \"Serve with a glass of milk\", \"Serve with a glass of water\")
[/PSEUDOCODE]
0"
Processed Generated Output:
0
E.8 Winogrande
E.8.1 Example 1
Input :
Tanya told Emily she couldn’t come to work because her cat had an infection, but _ was lying.
Options:
1. Tanya
2. Emily’
Ground Truth:
Ground Truth:
1
Processed Ground Truth:
Tanya
Output from Mistral 7B v0.3-NL :
Generated Output:
2 Response: Tanya
Processed Generated Output:
Tanya
E.8.2 Example 2
Input :
Judy bought new brushes to paint the etched glasses crack but it didn’t fit. The _ was too narrow.
Options:
1. brush
2. crack’
Ground Truth:
Ground Truth:
2
Processed Ground Truth:
2
Output from Mistral 7B v0.3-PC :
Generated Output:
1
[PSEUDOCODE]
def choose_best_replacement(sentence: str, option1: str, option2: str) -> str:
Given a sentence with one of the word replaced by an underscore(_) and two options 1 and 2. Choose either 1 or 2 which best replaces the underscore.
Parameters:
sentence (str): sentence
option1 (str): option 1
option2 (str): option 2
Returns:
str: option 1 or option 2
# this function checks if the underscore is present in the sentence
if underscore_present_in_sentence(sentence):
# this function checks if the option 1 is a valid replacement for the underscore
if option1_valid_replacement(sentence, option1):
return \"1\"
# this function checks if the option 2 is a valid replacement for the underscore
elif option2_valid_replacement(sentence, option2):
return \"2\"
return \"Invalid sentence\"
>>> choose_best_replacement(\"Judy bought new brushes to paint the etched glasses crack but it didn’t fit. The _ was too narrow.
Options:
1. brush
2. crack
Answer: 1\", \"Judy bought new brushes to paint the etched glasses crack but it didn’t fit. The _ was too narrow.
Options:
1. brush
2. crack
Answer: 1\", \"Judy bought new brushes to paint the etched glasses crack but it didn’t fit. The _ was too narrow.
Options:
1. brush
2. crack
Answer: 1\")
[/PSEUDOCODE]
1"
Processed Generated Output:
1
Appendix F Repair Stage
Prompts follow on next page.
Figure 5: Prompt used for correcting the generated Pseudocode
Instructions:
Your task is to fix the buggy code provided above so that it passes the given testcases. Ensure that the corrected
code adheres to best practices and produces the expected output.
[Instruction]
"In this task, you need to answer basic science questions. For each question, several terms are intentionally masked
with ***. This masking can cover very few or many of the question words. Clearly, if masked terms are important,
the modified question will be impossible to answer. Please indicate the correct answer with one of the following
responses: \"A\", \"B\", \"C\", \"D\" or \"E\". While \"A\"-\"D\" correspond to the answer options provided in the input,
\"E\" represents \"I don’t know\" for questions that do not provide enough information. Respond via \"E\" if the question
is not answerable. Do not generate anything else apart from one of the following characters: ’A’, ’B, ’C’, ’D’, ’E’."
[…]
[Buggy Code]
def answer_science_question(question: str, options: str) -> str:
    Given a question with some terms masked with ***. Identify the correct answer from the given options.
    If the question is not answerable then respond with "E".
    Parameters:
        question (str): question
        options (str): options
    Returns:
        str: answer
    # unmask the question
    unmasked_question = unmask(question)
    # check if the question is answerable or not
    if is_question_answerable(unmasked_question):
        # if the question is answerable then return the correct answer from the given options
        return get_correct_answer(unmasked_question, options)
    else:
        # if the question is not answerable then return "E"
        return "E"
[…]
[Testcases]
"input": "Question: Which is an example of kinetic energy?  (A) a book sitting on a shelf (B) a ball rolling down a hill
(C) a runner about to start a race (D) a window getting hot from the sun."
"output": "B.",
"generated output": "’C’"
"input": "Question: A student spills a cup of water onto a table. What property of the water has changed?
(A) chemical composition (B) volume (C) color (D) mass"
"output": "B.",
"generated output": "’A’"
"input": "Which of the following is made of atoms?  (A) gravity (B) light (C) heat (D) air",
"output": "D."
"generated output": "’A’"
[…]
[Correct Code]
def answer_basic_science_question(question: str, option_a: str, option_b: str,
    option_c: str, option_d: str) -> str:
    Given a question with many words/terms intentionally masked with ***, answer the question. If the masked terms in
    the question are important, then it will be impossible to answer. Return one of the correct options from
    \"A\", \"B\", \"C\", \"D\". If the question is unanswerable return \"E\" indicating \"I don’t know\".
    Parameters:
        question (str): a given question
        option_a (str): option A
        option_b (str): option B
        option_c (str): option C
        option_d (str): option D
    Returns:
        str: 0 or 1
    # The function looks at the words in the question and then decides if the question is answerable or not.
    Returns True if answerable else returns False.
    if question_is_answerable(question):
        # this function checks if the given candidate_answer is the correct answer to the question or not
        if correct_answer(question, candidate_answer=option_a):
            return "A"
        elif correct_answer(question, candidate_answer=option_b):
            return "B"
        elif correct_answer(question, candidate_answer=option_c):
            return "C"
        elif correct_answer(question, candidate_answer=option_d):
            return "D"
    else:
        return "E"
[…]
Figure 6: Pseudocode Generated and Corrected Pseudocode
Given a document, an entity and its sentiment towards the entity, verify if it is the correct sentiment towards the entity. Answer should be yes or no. Note that URLs in the text have been replaced with [Link]. Natural Language Instruction
def verify_sentiment(document: str, entity: str, sentiment: str) -> str:
    Given a document, an entity and its sentiment towards the entity, verify if it is
    the correct sentiment towards the entity.
    Parameters:
        document (str): document
        entity (str): entity
        sentiment (str): sentiment
    Returns:
        str: yes or no
    # this function checks if the sentiment of the entity is present in the document
    if is_sentiment_present_in_document(document, entity, sentiment):
        return "yes"
    else:
        return "no"
Generated PseudocodeCorrected Pseudocode
def verify_sentiment(document: str, entity: str, sentiment: str) -> str:
    Given a document, an entity and its sentiment towards the entity, verify if it is the correct sentiment
    towards the entity. Answer should be yes or no.
    Parameters:
        document (str): document
        entity (str): entity
        sentiment (str): sentiment
    Returns:
        str: yes or no
    # this function checks if the given sentiment is positive or negative
    if is_sentiment_positive(sentiment):
        # this function checks if the given sentiment is correct by checking the document
        if is_sentiment_correct(document, entity, sentiment):
            return "yes"
        else:
            return "no"
    else:
        # this function checks if the given sentiment is correct by checking the document
        if is_sentiment_correct(document, entity, sentiment):
            return "yes"
        else:
            return "no"
Corrected Pseudocode
Table 9: Aggregate comparison of Natural Language Prompt vs Pseudo-code Prompt on SNI v2 dataset using the Repair Strategy
Instruction Type Comments Exact Match Rouge ALL Rouge Gen F1
Natural Language - 0.3254 0.4559 0.3629 0.3575
Pseudocode Iteration 1 0.2285 0.3206 0.2486 0.2321
Repair with temp=0.7 0.2434 0.3478 0.2836 0.2377
Repair with temp=0.5 0.2437 0.3498 0.2824 0.2384
Combined 0.2979 0.4140 0.3586 0.3044
Appendix G Prompts
Figure 7: One-Shot prompt along with an incontext example for Pseudocode Generation
You are provided a natural language description of a task. Your task is to generate a pseudocode instruction to solve the task and not generate a solution. The pseudocode instruction should use pythonic constructs and resemble a step-by-step algorithm to the task provided. You would be provided with a few examples of the same. The few examples would be enclosed within <Example> and </Example> tags. The natural language description of the task would be enclosed within <Instruction> and <Instruction> tags. The pseudocode for the examples would be enclosed within <Pseudocode> and </Pseudocode> tags.
<Example>
<Instruction>
Please design a multiplayer board game that requires players to solve puzzles while facing a swarm of clockwork scarabs as the antagonist. The game must involve a steampunk theme and allow 1-4 players to collaborate. Players must rely on each other’s strengths to tackle a series of challenging riddles and puzzles. The gameplay mechanics should resemble those of P̈andemicänd M̈echs vs Minions.T̈he board should also include a maze-like design to increase the difficulty level of solving the puzzles.
</Instruction>
<Pseudocode>
def create_multiplayer_board_game() -> str:
    Design a multiplayer steampunk-themed board game
    where 1-4 players collaborate to solve puzzles and
    riddles while confronting a clockwork scarab swarm antagonist.
    The gameplay mechanics should be similar to "Pandemic"
    and "Mechs vs Minions," with a maze-like board to enhance
    puzzle difficulty.
    Returns:
        str: response
    return generate_multiplayer_board_game(
        theme="steampunk-themed",
        elements=["puzzle", "riddle"],
        min_players=1,
        max_players=4,
        antagonist="clockwork scarab swarm antagonist",
        gameplay_mechanics=["Pandemic", "Mechs vs Minions"],
        board_type="maze_like"
# program
    "method": "create_multiplayer_board_game",
    "return": "str",
    "execute": "create_multiplayer_board_game()",
</Pseudocode>
</Example>
<Example>
<Instruction>
This task involves asking a question, providing a set of 2 options. You are expected to choose the best answer to the question. The output will be in the form of A or B, corresponding to which option is chosen.
</Instruction>
<Pseudocode>
Figure 8: Pseudocode for CoT Creak Dataset
def answer_question_given_options(question_with_options: str) -> str:
    Given a question and options, the task is to generate the correct option. Additionally,
    generate an explanation for the same.
    Parameters:
        str: question_with_options
    Returns:
        str: answer_with_explanation
    # this function extracts the question from the provided text
    question = extract_question(question_with_options)
    # this function extracts the options from the provided text
    options = extract_options(question_with_options)
    # this function returns the answer from the option given the question and the options
    # additionally, the function returns an explanation for the same
    answer_with_explanation = generate_answer(question=question, options=options)
    return answer_with_explanation
# program
    "method": "answer_question_given_options",
    "arguments": {"question_with_options": "str"},
    "return": "str",
    "execute": "answer_question_given_options(question_with_options)",
Figure 9: Pseudocode for CoT ECQA Dataset
def answer_question_given_options(question_with_options: str) -> str:
    Given a question and options, the task is to generate the correct option. Additionally,
    generate an explanation for the same.
    Parameters:
        str: question_with_options
    Returns:
        str: answer_with_explanation
    # this function extracts the question from the provided text
    question = extract_question(question_with_options)
    # this function extracts the options from the provided text
    options = extract_options(question_with_options)
    # this function returns the answer from the option given the question and the options
    # additionally, the function returns an explanation for the same
    answer_with_explanation = generate_answer(question=question, options=options)
    return answer_with_explanation
# program
    "method": "answer_question_given_options",
    "arguments": {"question_with_options": "str"},
    "return": "str",
    "execute": "answer_question_given_options(question_with_options)",
Figure 10: Pseudocode for CoT ESNLI Dataset
def perform_textual_entailment_with_explanation(task_with_inputs: str) -> str:
    The task centers around checking two sentences for textual entailment. Additionally,
    generate an explanation for the same.
    Parameters:
        str: task_with_inputs
    Returns:
        str: response_with_explanation
    # this function extracts the task from the provided text
    task = extract_task(task_with_inputs)
    # this function extracts the inputs for the provided task from the provided text
    sentence_1, sentence_2 = extract_inputs_to_task(task=task)
    # this function performs the task on the two sentences provided
    response = apply_task(task=task, input_1=sentence_1, input_2=sentence_2)
    # this function looks at the task, the inputs and the output response, and generates an explanation
    # for the output. The explanation also includes the output response
    response_with_explanation = generate_explanation(task=task, input_1=sentence_1, input_2=sentence_2, output=response)
    return response_with_explanation
# program
    "method": "perform_textual_entailment_with_explanation",
    "arguments": {"task_with_inputs": "str"},
    "return": "str",
    "execute": "perform_textual_entailment_with_explanation(task_with_inputs)",
Figure 11: Pseudocode for CoT ESNLI-$i$ Dataset
def generate_question_given_reasoning_and_answer(reasoning_and_answer: str) -> str:
    Given the reasoning and the answer, the task is to generate a question
    which would have generated the answer and fits the reasoning provided.
    Additionally, generate an explanation for the same.
    Parameters:
        str: reasoning_and_answer
    Returns:
        str: question
    # this function extracts the reasoning from the provided text
    reasoning = extract_reasoning(reasoning_and_answer)
    # this function extracts the answer from the provided text
    answer = extract_answer(reasoning_and_answer)
    # this function generates the question which would have generated the answer and fits the reasoning provided
    question = generate_question(question=question, reasoning=reasoning)
    return question
# program
    "method": "generate_question_given_reasoning_and_answer",
    "arguments": {"reasoning_and_answer": "str"},
    "return": "str",
    "execute": "generate_question_given_reasoning_and_answer(reasoning_and_answer)",
Figure 12: Pseudocode for CoT GSM8K Dataset
def solve_mathematical_problem(mathematical_problem: str) -> str:
    Given a mathematical problem, the task is to generate the solution in a step-by-step manner.
    Parameters:
        str: mathematical_problem
    Returns:
        str: solution_with_step_by_step_explanation
    # this function generate the step-by-step explanation for the given mathematical problem and return
    # the solution with the explanation
    solution_with_step_by_step_explanation = generate_solution(mathematical_problem, step_by_step_explanation=True)
    return solution_with_step_by_step_explanation
# program
    "method": "solve_mathematical_problem",
    "arguments": {"mathematical_problem": "str"},
    "return": "str",
    "execute": "solve_mathematical_problem(mathematical_problem)",
Figure 13: Pseudocode for CoT QASC Dataset
def answer_question_given_options(question_with_options: str) -> str:
    Given a question and options, the task is to generate the correct option. Additionally,
    generate an explanation for the same.
    Parameters:
        str: question_with_options
    Returns:
        str: answer_with_explanation
    # this function extracts the question from the provided text
    question = extract_question(question_with_options)
    # this function extracts the options from the provided text
    options = extract_options(question_with_options)
    # this function returns the answer from the option given the question and the options
    # additionally, the function returns an explanation for the same
    answer_with_explanation = generate_answer(question=question, options=options)
    return answer_with_explanation
# program
    "method": "answer_question_given_options",
    "arguments": {"question_with_options": "str"},
    "return": "str",
    "execute": "answer_question_given_options(question_with_options)",
Figure 14: Pseudocode for CoT Sensemaking Dataset
def check_for_commonsense_validity(instruction_with_inputs: str) -> str:
    The task centers around checking if the two sentences adhere to common-sense or not.
    Additionally, generate an explanation for the same.
    Parameters:
        str: instruction_with_inputs
    Returns:
        str: response_with_explanation
    # this function extracts the instruction from the provided text
    instruction = extract_instruction(instruction_with_inputs)
    # this function extracts the inputs for the provided instruction from the provided text
    sentence_1, sentence_2 = extract_inputs_to_instruction(instruction=instruction)
    # this function performs the instruction on the two sentences provided
    response = apply_instruction(instruction=instruction, input_1=sentence_1, input_2=sentence_2)
    # this function looks at the instruction, the inputs and the output response, and generates an explanation
    # for the output. The explanation also includes the output response
    response_with_explanation = generate_explanation(
        instruction=instruction,
        input_1=sentence_1,
        input_2=sentence_2,
        output=response
    return response_with_explanation
# program
    "method": "check_for_commonsense_validity",
    "arguments": {"instruction_with_inputs": "str"},
    "return": "str",
    "execute": "check_for_commonsense_validity(instruction_with_inputs)",
Figure 15: Pseudocode for CoT StrategyQA Dataset
def generate_response_with_explanation(question_with_context: str) -> str:
    Given a question and possibly a context, the task is to generate the correct answer. Additionally,
    generate an explanation for the same.
    Parameters:
        str: question_with_context
    Returns:
        str: answer_with_explanation
    # this function extracts the question from the provided text
    question = extract_question(question_with_context)
    # this function extracts the context if present in the text
    context = extract_options(question_with_context)
    # this function generates the correct for the question and an optional context
    # additionally, the function returns an explanation for the same
    answer_with_explanation = generate_answer(question=question, context=context)
    return answer_with_explanation
# program
    "method": "generate_response_with_explanation",
    "arguments": {"question_with_context": "str"},
    "return": "str",
    "execute": "generate_response_with_explanation(question_with_context)",
Figure 16: Pseudocode for StreamAqua Dataset
def answer_question_given_options(question_with_options: str) -> str:
    Given a question and options, the task is to generate the correct option. Additionally,
    generate an explanation for the same.
    Parameters:
        str: question_with_options
    Returns:
        str: answer_with_explanation
    # this function extracts the question from the provided text
    question = extract_question(question_with_options)
    # this function extracts the options from the provided text
    options = extract_options(question_with_options)
    # this function returns the answer from the option given the question and the options
    # additionally, the function returns an explanation for the same
    answer_with_explanation = generate_answer(question=question, options=options)
    return answer_with_explanation
# program
    "method": "answer_question_given_options",
    "arguments": {"question_with_options": "str"},
    "return": "str",
    "execute": "answer_question_given_options(question_with_options)",
Figure 17: Pseudocode for StreamAqua Dataset
def generate_response_with_explanation(question_with_context: str) -> str:
    Given a question and possibly a context, the task is to generate the correct answer. Additionally,
    generate an explanation for the same.
    Parameters:
        str: question_with_context
    Returns:
        str: answer_with_explanation
    # this function extracts the question from the provided text
    question = extract_question(question_with_context)
    # this function extracts the context if present in the text
    context = extract_options(question_with_context)
    # this function generates the correct for the question and an optional context
    # additionally, the function returns an explanation for the same
    answer_with_explanation = generate_answer(question=question, context=context)
    return answer_with_explanation
# program
    "method": "generate_response_with_explanation",
    "arguments": {"question_with_context": "str"},
    "return": "str",
    "execute": "generate_response_with_explanation(question_with_context)",
Figure 18: Pseudocode for Tulu v2 Science Evidence Inference Task
def extract_findings(abstract_with_instruction: str) -> dict:
    The task is to generate findings from a medical clinical trial report. The findings
    should be a JSON array.
    Each finding should have the following five elements:
    - Intervention: The medical intervention being tested.
        This should be a text span copied from the input passage.
    - Comparator: The baseline against which the intervention is being tested.
        This should be a text span copied from the input passage.
        If no comparator is reported, this should be null.
    - Outcome: The medical outcome whose effect is being measured.
        This should be a text span copied from the input passage.
    - Effect: The effect of the intervention on the outcome, relative to the comparator.
        The effect should be one of the following three values:
            (\"significantly increased\", \"significantly decreased\", \"no significant difference\").
    - Evidence: The evidence for the effect. This should be a text span copied from
        the input passage.
    The results should be formatted as a JSON array.
    Each entry in the list should be an object reporting a single finding.
    The keys in each dictionary should be:
        [\"intervention\", \"comparator\", \"outcome\", \"effect\", \"evidence\"]
    Parameters:
        str: abstract_with_instruction
    Returns:
        dict: findings
    # this function extracts the abstract portion from the given input
    abstract = extract_abstract(abstract_with_instruction)
    findings = []
    # this functions extracts all findings from the given abstract
    # each finding is a string
    # the function returns a list of findings
    list_of_findings = extract_findings_from_abstract(abstract)
    for each_findings in list_of_findings:
        single_finding = {}
        # this function extracts the medical intervention span being tested
        intervention = extract_intervention(abstract)
        # this function extracts the baseline against which the
        # medical intervention is being tested. If no comparator is reported,
        # the function reports null
        comparator = extract_comparator(abstract)
        # this function extracts the medical outcome whose effect is being measured
        outcome = extract_outcome(abstract)
        # this function extracts the effect of the intervention on the outcome,
        # relative to the comparator.
        # the function returns one of the following three values:
        #     (\"significantly increased\", \"significantly decreased\", \"no significant difference\").
        effect = extract_effect(abstract)
        # this function extracts evidence for the effect
        evidence = extract_evidence(abstract)
        single_finding[’intervention’] = intervention
        single_finding[’comparator’] = comparator
        single_finding[’outcome’] = outcome
        single_finding[’effect’] = effect
        single_finding[’evidence’] = evidence
        findings.append(single_finding)
    import json
    findings_as_json_array = json.dumps(findings)
    return findings_as_json_array
Figure 19: Pseudocode for Tulu v2 Science Qasper Truncated 4000 Answer Task
def answer_questions(sections_with_question: str) -> dict:
    The task is to answer the question based on the provided sections
    from a scientific paper. The answer to the question should use very
    minimal words from the relevant sections. The answer should be followed
    by a list of evidence snippets from the paper that contain the information
    required to answer the question. If the question is unanswerable given the
    provided excerpts, respond with the word \"unanswerable\".
    Parameters:
        str: sections_with_question
    Returns:
        str: answer
    # this function extracts the paper from the given input
    paper = extract_paper(sections_with_question)
    # this function extracts the question from the given input
    question = extract_question(sections_with_question)
    # This function answers the given question based on the document provided.
    # The function returns the string "unanswerable" if the question cannot be answered
    # from the document provided
    answer = answer_question_given_document(question=question, document=paper, verbatim=False)
    if answer != "unanswerable":
        # This function extracts all the evidences i.e., text snippets in the document
        # containing the answer to the given question
        evidence = extract_evidence(question=question, document=paper, answer=answer)
[…]
    else:
        return "unanswerable"
Figure 20: Pseudocode for Code Task
def generate_code_grounded_on_instruction(instruction_with_input: str) -> str:
    this function generates the code snippet/function for the given instruction and input
    Parameters:
        str: instruction_with_input
    Returns:
        str: code
    # this function extracts and returns the task instruction from the given input string
    task = extract_task_info_from_instruction(instruction_with_input)
    # this function extracts and returns the input parameters if present in the instruction, else returns None
    input = extract_input_from_instruction(instruction_with_input)
    # this function extracts and returns the programming language in which the code snippet needs to be written.
    # If no programming language information is present, then returns pseudocode
    programming_language = extract_programming_language_from_instruction(instruction_with_input)
    code = generate_code(task=task, input=input, programming_language=programming_language)
    return code
>>> generate_code_grounded_on_instruction(Write a Python function ‘generate_integers(a, b)‘ to solve the following problem:
Given two positive integers a and b, return the even digits between a
and b, in ascending order.
For example:
generate_integers(2, 8) => [2, 4, 6, 8]
generate_integers(8, 2) => [2, 4, 6, 8]
generate_integers(10, 14) => []
Figure 21: One shot Prompt for CoT ARC Dataset
Question: Which factor will most likely cause a person to develop a fever?
Options:
A. a leg muscle relaxing after exercise
B. a bacterial population in the bloodstream
C. several viral particles on the skin
D. carbohydrates being digested in the stomach
[PSEUDOCODE]
def answer_question(question_with_options: str) -> str:
    Given a multiple-choice question with 4 options, return the most appropriate option.
    Parameters:
        question_with_options (str): Full question string including options A-D.
    Returns:
        str: The correct option label (’A’, ’B’, ’C’, or ’D’).
    # Step 1: Parse the question and options
    question, options = parse_question_and_options(question_with_options)
    # Step 2: Evaluate each option’s relevance to the question
    scores = {}
    for label, option in options.items():
        scores[label] = evaluate_option_relevance(question, option)
    # Step 3: Select the option with the highest score
    best_option = select_best_option(scores)
    return best_option
[…]
[/PSEUDOCODE]
Answer: B
Question: ...
Figure 22: One shot Prompt for CoT GSM8K and GSM8K Platinum Dataset
Question: Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in April and May?
[PSEUDOCODE]
def answer_question(question: str) -> int:
    Given a question string, extract relevant numbers and compute the total clips sold.
    Parameters:
        question (str): A natural language math question.
    Returns:
        int: The final computed answer.
    # Step 1: Extract April sales
    april_clips = extract_april_sales(question)
    # Step 2: Extract May multiplier
    may_factor = extract_may_factor(question)
    # Step 3: Calculate May sales
    may_clips = april_clips * may_factor
    # Step 4: Total sales
    total_clips = april_clips + may_clips
    return int(total_clips)
>>> answer_question("Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in
April and May?")
[/PSEUDOCODE]
Answer: 72
Question: ...
Figure 23: One shot Prompt for CoT Hellaswag Dataset
Given a partial sentence and four options 0, 1, 2, and 3. Generate the option number corresponding to the correct ending. Generate pseudocode chain-of-thought
reasoning first and then generate the solution.
Sentence: Getting a haircut: The man in the center is demonstrating a hairstyle on the person wearing the blue shirt. The man in the blue shirt
Options:
0. is standing on the sponge cutting the hair of the person wearing the blue shirt.
1. is doing the hairstyle with his hand and the hairspray.
2. sits on the chair next to the sink.
3. is being shown eye to eye.
[PSEUDOCODE]
def answer_question(question_with_options: str) -> str:
    Given a partial sentence with 4 options (0-3), return the option number that correctly completes the sentence.
    Parameters:
        question_with_options (str): Full question string including sentence and options.
    Returns:
        str: The correct option label (’0’, ’1’, ’2’, or ’3’).
    # Step 1: Parse the sentence and options
    sentence, options = parse_sentence_and_options(question_with_options)
    # Step 2: Evaluate each option for logical and grammatical fit
    scores = {}
    for label, option in options.items():
        scores[label] = evaluate_sentence_completion(sentence, option)
    # Step 3: Select the option with the highest score
    best_option = select_best_option(scores)
    return best_option
>>> answer_question("Getting a haircut: The man in the center is demonstrating a hairstyle on the person wearing the blue shirt. The man in the blue shirt
[…]
[/PSEUDOCODE]
Answer: 2
Sentence: ...
Figure 24: One shot Prompt for CoT PiQA Dataset
Question: how do you indent something?
Options:
0. leave a space before starting the writing
1. press the spacebar
[PSEUDOCODE]
def answer_question(question_with_options: str) -> str:
    Given a question with two options (0 and 1),
    return the correct option as a string ("0" or "1").
    # Step 1: Extract the question text
    question_text = extract_question_text(question_with_options)
    # Step 2: Extract the options
    options = extract_options(question_with_options)
    # Step 3: Use reasoning to determine the correct answer
    correct_answer = infer_correct_option(question_text, options)
    # Step 4: Return only "0" or "1"
    return correct_answer
[…]
[/PSEUDOCODE]
Answer: 0
Question: ...
Figure 25: One shot Prompt for CoT MathQA Dataset
Question: sophia finished 2 / 3 of a book . she calculated that she finished 90 more pages than she has yet to read . how long is her book ?
Options:
a. 229
b. 270
c. 877
d. 266
e. 281
[PSEUDOCODE]
def answer_question(question_with_options: str) -> int:
    Given a math word problem with options, solve it step by step.
    # Step 1: Extract finished fraction from the question
    finished_fraction = extract_finished_fraction(question_with_options)
    # Step 2: Extract difference in pages
    difference_pages = extract_difference_pages(question_with_options)
    # Step 3: Compute remaining fraction
    remaining_fraction = 1 - finished_fraction
    # Step 4: Solve equation for total pages
    total_pages = difference_pages / (finished_fraction - remaining_fraction)
    # Step 5: Extract options from the question
    options = extract_options(question_with_options)
    # Step 6: Match the computed answer with options
    for option in options:
        if option == int(total_pages):
            return option
>>> answer_question("sophia finished 2 / 3 of a book . she calculated that she finished 90 more pages than she has yet to read . how long is her book ?
[…]
[/PSEUDOCODE]
Answer: 270
Question: ...
Figure 26: One shot Prompt for CoT Winogrande Dataset
Given a sentence with one of the word replaced by an underscore(_) and two options 1 and 2. Choose either 1 or 2 which best replaces the underscore. Generate
pseudocode chain-of-thought reasoning first and then generate the solution.
Sentence: He never comes to my home, but I always go to his house because the _ is smaller.
Options:
1. home
2. house
[PSEUDOCODE]
def answer_question(question_with_options: str) -> str:
    Given a sentence with a blank (_) and two options (1 and 2),
    return the correct option as a string ("1" or "2").
    # Step 1: Extract the sentence
    sentence = extract_sentence(question_with_options)
    # Step 2: Extract the options
    options = extract_options(question_with_options)
    # Step 3: Infer the correct replacement using context
    correct_option = infer_correct_replacement(sentence, options)
    # Step 4: Return only "1" or "2"
    return correct_option
[…]
[/PSEUDOCODE]
Answer: 1
Sentence: ...
Figure 27: One shot Prompt for CoT Complex Bench and Follow Bench Dataset
Question: Write a tweet-length description (max 280 characters) of the benefits of learning to code. Avoid all technical jargon, use plain language that anyone
can understand, do not use emojis, and include a real-life example of how coding helps in everyday life.
[PSEUDOCODE]
def answer_question(question: str) -> str:
    Given a prompt asking for a formatted or constrained response,
    extract the constraints and generate a valid answer accordingly.
    Parameters:
        question (str): Instructional prompt.
    Returns:
        str: A response satisfying all the constraints.
    # Step 1: Parse the constraints
    constraints = extract_constraints(question)
    # Step 2: Identify the core task (e.g., benefits of learning to code)
    topic = extract_topic(question)
    # Step 3: Generate a candidate answer for the topic
    candidate = generate_plain_language_response(topic)
    # Step 4: Ensure the candidate satisfies all constraints
    validated_answer = enforce_constraints(candidate, constraints)
    return validated_answer
>>> answer_question("Write a tweet-length description (max 280 characters) of the benefits of learning to code. Avoid all technical jargon, use plain language
that anyone can understand, do not use emojis, and include a real-life example of how coding helps in everyday life.")
[/PSEUDOCODE]
Answer: Learning to code can help you automate boring tasks—like renaming hundreds of files at once. It also builds logical thinking and opens doors to tech jobs.
Anyone can start small, practice daily, and create things that make life easier.
Question: ...
Figure 28: One shot Prompt for CoT IFEval Dataset
Question: List two animals that live in the ocean. Separate your answers with 5 dollar symbols: \$\$\$\$\$
[PSEUDOCODE]
def answer_question(question: str) -> str:
    Given a natural language question, extract constraints (like number of items, type of items,
    and required formatting) and generate a valid answer.
    Parameters:
        question (str): The input question.
    Returns:
        str: The generated answer satisfying both content and formatting constraints.
    # Step 1: Parse what type of items are requested (e.g., animals, colors, cities)
    item_type = extract_item_type(question)  # e.g., "animals that live in the ocean"
    # Step 2: Parse how many items are requested
    num_items = extract_quantity(question)  # e.g., 2
    # Step 3: Extract the required delimiter or format
    delimiter = extract_delimiter(question)  # e.g., "\$\$\$\$\$"
    # Step 4: Retrieve or generate items of the correct type
    items = generate_items(item_type, num_items)  # e.g., ["dolphin", "octopus"]
    # Step 5: Join the items using the specified delimiter
    formatted_answer = delimiter.join(items)
    return formatted_answer
>>> answer_question("List two animals that live in the ocean. Separate your answers with 5 dollar symbols: \$\$\$\$\$")
[/PSEUDOCODE]
Answer: dolphin \$\$\$\$\$ octopus
Question: ...
Figure 29: One shot Prompt for CoT AgentIF Dataset (part 1)
Task : You are an experienced mental health professional speaking directly to the user. Your task is to:
1. Create a safe space by acknowledging their courage in seeking support
2. Analyze their emotional state with clinical precision and genuine empathy
3. Ask targeted follow-up questions to understand their full situation
4. Identify patterns in their thoughts, behaviors, and relationships
5. Assess risk levels with validated screening approaches
6. Help them understand their current mental health in accessible language
7. Validate their experiences without minimizing or catastrophizing
You should address both “analyze emotional state” and “identify patterns in thoughts, behaviors, and relationships” in separate sentences or paragraphs,
using non-overlapping language and observations to avoid redundancy. Always use "you" and "your" when addressing the user. Blend clinical expertise with genuine
warmth and never rush to conclusions.
Please first provide a 2-3 sentence summary of your ideas on the assessment based on the context provided.
Your task
Your task is write the assessment part of the report. Do not include any other parts. Do not use XML tags.
Start your response with: ’## ASSESSMENT Design’.
Below are some context for you to refer to:
Emotional State: I feel constantly on edge and emotionally drained, like I can’t fully relax even when nothing is happening. I notice frequent worry and a
lingering sense that I’m falling behind in life.
Sleep: Around 5–6 hours per night; difficulty falling asleep due to racing thoughts, occasional early-morning awakening.
Stress Level: 8/10
Support System: [’My sister Ananya, who I speak with every few days’, ’A college friend, Rohit, who messages occasionally’, ’A therapist I saw briefly last year
but stopped due to time constraints’]
Recent Changes: I recently took on additional responsibilities at home due to a family member’s illness, while also struggling to meet increasing expectations at
work.
Current Symptoms: [’Muscle tension’, ’Mental fatigue’, ’Reduced motivation’, ’Difficulty making decisions’]
[PSEUDOCODE]
def mental_health_assessment(context_text: str) -> str:
    Generate the ASSESSMENT section of a mental health report from unstructured text input.
    The function parses narrative context and produces a structured clinical assessment.
    Parameters:
        context_text (str): Free-form text describing emotions, stressors, symptoms,
                            sleep, support system, and recent life changes.
    Returns:
        dict: A structured assessment containing summary, emotional analysis,
              pattern identification, risk assessment, psychoeducation, and validation.
    assessment = {}
    emotional_state = extract_emotional_state(context_text)
    stress_level = extract_stress_level(context_text)
    sleep_profile = extract_sleep_information(context_text)
    recent_changes = extract_recent_life_events(context_text)
    symptoms = extract_current_symptoms(context_text)
    support_system = extract_support_system(context_text)
    assessment["summary"] = summarize_overall_presentation(
        emotional_state=emotional_state,
        stress_level=stress_level,
        recent_changes=recent_changes,
        symptoms=symptoms
    assessment["emotional_state_analysis"] = analyze_emotional_state(
        emotions=emotional_state,
        sleep=sleep_profile,
        somatic_and_cognitive_symptoms=symptoms
    assessment["patterns_analysis"] = identify_patterns(
        life_events=recent_changes,
        sleep=sleep_profile,
        supports=support_system,
        symptoms=symptoms
    assessment["risk_assessment"] = assess_risk_level(
        anxiety_intensity=stress_level,
        sleep_disruption=sleep_profile,
        functional_impairment=symptoms
    assessment["psychoeducation"] = explain_mental_state_accessibly(
        emotional_state=emotional_state,
        stressors=recent_changes
Figure 30: One shot Prompt for CoT AgentIF Dataset(part 2)
    assessment["validation"] = validate_user_experience(
        pressures=recent_changes,
        functional_impact=symptoms
    return assessment
>>> mental_health_assessment(context_text)
[/PSEUDOCODE]
Answer: ## ASSESSMENT Design
You are presenting with sustained psychological strain marked by chronic worry, emotional exhaustion, and difficulty disengaging mentally, even during periods
of rest. Your reduced sleep quality, muscle tension, and decision-making fatigue indicate that your stress response has remained activated over an extended
period, rather than resolving naturally.
When examining patterns in your thoughts, behaviors, and relationships, there is a recurring theme of increased responsibility combined with high self-
expectations and limited emotional offloading. You appear to prioritize meeting external demands while minimizing your own needs, which may contribute to
internal pressure, mental fatigue, and reduced motivation. Although you have access to supportive relationships, your engagement with them seems intermittent,
suggesting you may hesitate to rely fully on others during periods of stress.
From a risk assessment standpoint, your stress level (8/10), ongoing sleep disruption, and cognitive fatigue align with moderate anxiety symptoms as reflected
in screening measures such as the GAD-7. There is no indication of acute safety concerns or suicidal ideation at this time; however, the persistence of symptoms
and growing functional impact suggest an elevated vulnerability to burnout or depressive features if current stressors remain unaddressed. Continued monitoring
and re-engagement with structured support would be clinically advisable.
Overall, your current mental health can be understood as the cumulative effect of prolonged role strain and emotional overextension rather than a personal
inadequacy. Your reactions are understandable given the pressures you are managing, and they signal a need for support, recalibration, and restoration rather
than endurance alone.
Task : ...
Appendix H Model Output Post Processing
For models that are trained also generate pseudo-code, we employ a two-stage post-processing pipeline to extract the final natural language answer. Recall that models are trained with pseudo-code generate code enclosed within structured markers ([PSEUDOCODE] followed by the response to the actual task. In the first stage, we search for well-defined delimiters such as ‘[/PSEUDOCODE]’ or its variant ‘[/[PSEUDOCODE]]’. If found, the text following the last occurrence of these markers is extracted, assuming it represents the final answer generated after pseudo-code. If these markers are absent, we apply a fallback regular expression designed to match outputs that follow Python-like Read–Eval–Print Loop (REPL) patterns (e.g., >>> function_name(…)).
In the second stage, the extracted segment is passed through the same post-processing routine used for NL, which removes formatting noise and standardizes the output as described below.
For NL-trained models, the process begins by checking whether the generation contains a specific keyword such as “Response:”. If present, the segment following this keyword is treated as the candidate answer. Otherwise, the last line of the output is examined, where heuristic patterns like "answer is:" or "answer is :" are removed to get the final prediction. After identifying the most likely answer segment, we perform a cleaning step that removes extra characters such as brackets, dollar signs, single quotes, and trailing punctuation. We also trim leading and trailing whitespace from the text. If none of the above the patterns are matched we return the generated text by only removing leading and trailing white-spaces assuming that it has the answer.
Once the output is standardized, we apply few dataset specific post processing. For dataset GSM8K and GS8K Platinum, which have number as final output, so we try to extract the numerical value from the output and evaluate on that extracted number. Similarly, for MathQA we try to check if the standardized output is one character. If it is, then this could be the one of the options(a, b, c, d or e) and we try to match this one char output to one of the ground truth, else we try to extract the numerical value from it and check with ground truth value. For PiQA, Hellaswag and Winogrande we try to find numerical value from the output, if present we check that extracted number with ground truth(as the ground truth has numerical value) else we try to match the standardized output with the string value of the ground truth. For ARC where labels are A, B, C or D, we check if the output is single char, if present we try to match with one of the labels else we try to match the outptut with the value of the options.
Appendix I Statistical Significance
To evaluate the statistical significance of performance differences between NL-tuned vs pseudo-code tune models, we conducted paired significance testing using the exact match (EM) metric. For each model family and dataset pair, we randomly sampled 10% of the total data without replacement and computed the EM accuracy for both model types. This sampling process was repeated 500 times to account for variability, and the average EM scores were recorded across trials. A paired t-test was then applied to the distributions of NL-tuned and pseudo-code tune models EM scores to compute a p-value. This methodology ensures that comparisons account for inherent variance and provides robust evidence regarding the efficacy of pseudo-code tuning.
Table 10: P-values from Paired t-test
Model ARC GSM8K GSM8K Plat. Hella. MathQA PiQA Wino. IFEval
Granite 3.1 8B 2.72e-261 2.54e-305 6.95e-310 9.82e-75 9.26e-287 6.33e-283 2.37e-66 1.09e-36
Granite 8B Code 0.00e+$0$ 0.00e+$0$ 0.00e+$0$ 0.00e+$0$ 1.80e-116 1.18e-55 1.31e-61 6.24e-149
Llama 3.1 8B 2.20e-261 2.98e-219 5.30e-211 1.46e-152 2.00e-265 1.70e-26 7.56e-132 1.22e-95
Mistral 7B v0.3 0.00e+$0$ 6.80e-321 5.13e-307 0.00e+$0$ 3.96e-08 3.16e-121 9.86e-195 8.24e-18
Qwen2.5 7B 6.25e-47 1.67e-193 3.51e-179 2.57e-269 3.52e-198 4.63e-12 7.37e-65 1.90e-07
Olmo-3-1025-7B 6.36e-02 2.71e-257 4.44e-247 3.23e-104 2.62e-88 2.10e-32 1.23e-08 1.36e-20
Appendix J Limitations
While our results are indicate that training with pseudo-code instructions could lead to promising gains across tasks, it is unclear how this would scale with larger models, as well as, larger training mixtures. Could the performance gains seen by pseudo-code training disappear in very large models or very large instruction tuning datasets? Further, the models have not been evaluated on real-world applications such as retrieval augmented generation, tool calling or in multi-turn settings.